### v2

# Robust Max Statistics for High-Dimensional Inference

Abstract
Although much progress has been made in the theory and application of bootstrap approximations for max statistics in high dimensions, the literature has largely been restricted to cases involving light-tailed data. To address this issue, we propose an approach to inference based on robust max statistics, and we show that their distributions can be accurately approximated via bootstrapping when the data are both high-dimensional and heavy-tailed. In particular, the data are assumed to satisfy an extended version of the well-established $L^{4}$-$L^{2}$ moment equivalence condition, as well as a weak variance decay condition. In this setting, we show that near-parametric rates of bootstrap approximation can be achieved in the Kolmogorov metric, independently of the data dimension. Moreover, this theoretical result is complemented by encouraging empirical results involving both Euclidean and functional data.
1 Introduction
Over the past decade, distributional approximation results for max statistics have become a prominent topic in high-dimensional inference. A prototypical example of such a statistic has the form $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$, where $\bar{X}\in\mathbb{R}^{p}$ is the sample mean vector of $n$ observations and $\mu=\mathbf{E}(\bar{X})$, but numerous variants arise in diverse contexts. Indeed, one of the main drivers of research on this topic is that many high-dimensional inference tasks can be unified within the problem of approximating the distribution of $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$, or some adaptation of it. For instance, such approximations can be directly applied to construct simultaneous tests and confidence intervals for coordinate-wise means $\mu_{1},\dots,\mu_{p}$. More broadly, other applications include detection of treatment effects (Sun:2022), error estimation for sample covariance matrices (Lopes:Bernoulli), post-selection inference (Kuchibhotla:2020:PSI), change-point detection (Yu:2021), confidence bands in non-parametric regression (Singh:2023), tests for shape restrictions (Chetverikov:shape), and more. Meanwhile, another major reason why max statistics have attracted growing interest is that bootstrap methods can accurately approximate their distributions when $p$ is much larger than the sample size $n$, which has been demonstrated by a cascade of theoretical advances CCK:2013; CCK:2017; Deng:2020; Kuchibhotla2020high; lopes2020bootstrapping; Kuchibhota_2021; Lopes:2022:AOS; CCK:2023; Fang:2023; Koike:2024.
Despite the substantial innovations that have been made in bootstrap approximations for max statistics, there is an Achilles heel that continues to hinder much of the research in this area. Namely, there is a widespread reliance on the assumption that the covariates have light tails, e.g., sub-Gaussian or sub-exponential. Moreover, there are empirical and theoretical results suggesting that light tails are necessary for bootstrap methods to successfully approximate the distributions of conventional max statistics in high dimensions (zhang2017gaussian; Giessing:2020; Preinerstorfer:2024). For instance, the simulations in (Giessing:2020) show that the Gaussian multiplier bootstrap performs poorly for $\max_{1\leq j\leq p}\sqrt{n}|\bar{X}_{j}-\mu_{j}|$ when the covariates have heavy tails and $p\gg n$. From a theoretical standpoint, it has also been proven that there is a moment-dependent phase transition governing the success of Gaussian approximations for $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$ (zhang2017gaussian; Preinerstorfer:2024). That is, if $W\in\mathbb{R}^{p}$ is a centered Gaussian vector having the same covariance matrix as $\sqrt{n}(\bar{X}-\mu)$, then the Kolmogorov distance between $\max_{1\leq j\leq p}W_{j}$ and $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$ may or may not vanish in the limit that $n$ and $p$ jointly diverge, depending on whether the covariates have enough moments. This breakdown of Gaussian approximations suggests that similar behavior should occur for bootstrap approximations—especially in the case of the Gaussian multiplier bootstrap, which seeks to mimic the behavior of $\sqrt{n}(\bar{X}-\mu)$ by generating random vectors from a centered Gaussian distribution whose covariance matrix is an estimate for that of $\sqrt{n}(\bar{X}-\mu)$.
Due to the issues just mentioned, there are strong motivations to extend bootstrap methods involving max statistics so that they can be applied reliably to high-dimensional data with heavy tails. However, the research in this direction is still at a very early stage, and there are just a couple of previous works that have given it attention. The first of these works briefly outlined an approach that combines truncation with permutation-based sampling (Lou:2017), but it was ultimately not pursued as a practical method for inference. More recently, the state-of-the-art paper (Fan:2023) proposed a weighted bootstrap for a max statistic of the form $\max_{1\leq j\leq p}\sqrt{n}|\hat{\theta}_{j}-\theta_{j}|$, where $\theta_{j}$ denotes the so-called “pseudomedian” of the $j$th covariate, and $\hat{\theta}_{j}$ is the classical Hodges-Lehmann estimator for $\theta_{j}$ (Hodges:1963).
While the approach in (Fan:2023) achieved major progress by delivering robust simultaneous inference for $\theta_{1},\dots,\theta_{p}$, it still has some essential limitations. One is that the pseudomedians can be unsatisfactory substitutes for the means $\mu_{1},\dots,\mu_{p}$, particularly in cases of asymmetric distributions, for which $\theta_{j}$ and $\mu_{j}$ may be quite different. A related issue is that an approach based on pseudomedians does not extend naturally to suprema of zero-mean empirical processes, which appear frequently in applications of bootstrap approximations for max statistics (CCK_bands; Chen:ridges; Han:Bernoulli; Chen:Jackknife; Dette:Banach; Lopes:Bernoulli; Giessing:processes). Another issue is that the method in (Fan:2023) produces simultaneous confidence intervals for $\theta_{1},\dots,\theta_{p}$ that are only theoretically justified when they all have the same width, which is impractical if the covariates fluctuate over different scales. Lastly, the available theoretical analysis for $\max_{1\leq j\leq p}\sqrt{n}|\hat{\theta}_{j}-\theta_{j}|$ establishes a near $n^{-1/4}$ rate of bootstrap approximation in the Kolmogorov metric, which does not align with other recent results for max statistics that establish near $n^{-1/2}$ rates in the setting of light-tailed data lopes2020bootstrapping; Lopes:2022:AOS; CCK:2023; Fang:2023; Koike:2024.
In the current paper, we propose to bootstrap a robust max statistic that enables simultaneous inference on the means $\mu_{1},\dots,\mu_{p}$ and overcomes the difficulties described above. Our approach is designed in terms of three ingredients: truncation, partial standardization, and the median-of-means (mom) technique  (Nemirovsky1983; lugosi2019mean). To briefly lay out the main ideas, let $X_{1},\dots,X_{n}\in\mathbb{R}^{p}$ be i.i.d. observations with $\textup{var}(X_{1j})=\sigma_{j}^{2}$, and $\mu_{j}=\mathbf{E}(X_{1j})$ as before. Also, let $\hat{\sigma}_{1}^{2},\dots,\hat{\sigma}_{p}^{2}$ denote variance estimates that will be constructed from a small hold-out set via mom, and define the truncation function $\varphi_{t}(x)=\textup{sgn}(x)(|x|\wedge t)$ for any $x\in\mathbb{R}$ and $t\geq 0$, where $a\wedge b=\min\{a,b\}$. In this notation, the proposed robust max statistic is defined by
$\small\mathcal{M}_{n}=\max_{1\leq j\leq p}\sum_{i=1}^{n}\frac{\varphi_{\hat{t}_{j}}(X_{ij}-\mu_{j})}{\hat{\sigma}_{j}^{\tau}n^{1/2}},$ (1)
where $\hat{t}_{j}=\sqrt{n}\hat{\sigma}_{j}$ for $j=1,\dots,p$, and $\tau\in[0,1]$ is a fixed partial standardization parameter.
Importantly, there is a direct link between distributional approximations for $\mathcal{M}_{n}$ and inference on the means $\mu_{1},\dots,\mu_{p}$. This is due to the monotonicity of the functions $\varphi_{\hat{t}_{j}}(\cdot)$, which makes it straightforward to construct simultaneous confidence intervals for the means using quantile estimates for $\mathcal{M}_{n}$ and its corresponding min statistic, as discussed in Section 2. The robust variance estimates $\hat{\sigma}_{j}^{2}$ also play an essential role, because they ensure that the confidence intervals induced by $\mathcal{M}_{n}$ are automatically adapted to the scale of the covariates, which is an issue that has often been neglected in the literature on max statistics.
For the purpose of bootstrapping $\mathcal{M}_{n}$, let $\tilde{X}_{j}$ denote a hold-out mom estimate of $\mu_{j}$ to be defined later, and let $\bar{\varphi}_{j}=\frac{1}{n}\sum_{i=1}^{n}\varphi_{\hat{t}_{j}}(X_{ij}-\tilde{X}_{j})$. In addition, let $\xi_{1},\dots,\xi_{n}\sim N(0,1)$ be i.i.d. Gaussian multipliers generated independently of the data. Putting these pieces together, we define a bootstrap sample of $\mathcal{M}_{n}$ as
$\small\mathcal{M}_{n}^{*}=\max_{1\leq j\leq p}\sum_{i=1}^{n}\frac{\xi_{i}\big(\varphi_{\hat{t}_{j}}(X_{ij}-\tilde{X}_{j})-\bar{\varphi}_{j}\,\big)}{\hat{\sigma}_{j}^{\tau}n^{1/2}}.$ (2)
With regard to theoretical analysis, we focus on a setting where the tails of the data are quantified by a variant of the $L^{4}$-$L^{2}$ moment equivalence condition, which has gained increasing currency in the high-dimensional robustness literature lugosi2019mean; ke2019user; mendelson2020robust; roy2021empirical; abdalla2022covariance. Specifically, we assume there is some $\delta>0$ such that the bound $\|\langle v,X_{1}-\mu\rangle\|_{L^{4+\delta}}\lesssim\|\langle v,X_{1}-\mu\rangle\|_{L^{2}}$ holds for all $v\in\mathbb{R}^{p}$, and in Proposition 1, we show that this condition is satisfied by heavy-tailed instances of well-known models. The other primary structural assumption in our analysis is that the covariates have a weak variance decay property of the form $\sigma_{(j)}^{2}\asymp j^{-2\beta}$ for some fixed $\beta>0$, where $\sigma_{(1)}^{2}\geq\cdots\geq\sigma_{(p)}^{2}$ are the sorted coordinate-wise variances. Notably, the decay is referred to as weak because the parameter $\beta$ is allowed to be arbitrarily small. Furthermore, it is known that this type of structure arises naturally in a variety of high-dimensional contexts that are related to principal components analysis and functional data analysis, among others (lopes2020bootstrapping). Under the complete set of conditions given in Assumption 1, our main result shows that with high probability, the Kolmogorov distance $\sup_{s\in\mathbb{R}}|\mathbf{P}(\mathcal{M}_{n}\leq s)-\mathbf{P}(\mathcal{M}_{n}^{*}\leq s|X)|$ is nearly of order $n^{-1/2}$, where $\mathbf{P}(\cdot|X)$ denotes probability that is conditional on all of the observations.
From a practical standpoint, the proposed method has several strengths. First, the method does not require fine tuning, which is demonstrated by the fact that we use the simple choices of $\hat{t}_{j}=\sqrt{n}\hat{\sigma}_{j}$ and $\tau=0.9$ throughout all of the experiments presented in Section 4. Second, we show that the method reliably produces well-calibrated tests and confidence intervals across many conditions—including heavy-tailed data generated from separable and elliptical distributions, as well as heavy-tailed functional data with rough sample paths. Third, the simulation results reveal that the proposed method performs favorably in comparison to the pseudomedian approach in (Fan:2023).
Notation. If $A$ is a real matrix, its Frobenius norm is $\|A\|_{F}=\sqrt{\textup{tr}(A^{\top}A)}$, and its operator norm $\|A\|_{\textup{op}}$ is the same as its largest singular value. If $x$ and $y$ are Euclidean vectors of the same dimension, then $\langle x,y\rangle$ denotes the Euclidean inner product, and $\|x\|_{2}=\sqrt{\langle x,x\rangle}$. If $\xi$ is a scalar random variable and $1\leq q<\infty$, we write $\|\xi\|_{L^{q}}=(\mathbf{E}|\xi|^{q})^{1/q}$, and in the case when $q=\infty$, we use $\|\xi\|_{L^{\infty}}$ to refer to the essential supremum. If $f$ is a scalar-valued function on $\mathbb{R}$, the notation $\|f\|_{L^{\infty}}$ is understood analogously with respect to Lebesgue measure. If $\{a_{n}\}$ and $\{b_{n}\}$ are sequences of non-negative real numbers, then the relations $a_{n}\lesssim b_{n}$ and $a_{n}=\mathcal{O}(b_{n})$ are equivalent, and mean that there is a constant $c>0$ not depending on $n$, such that $a_{n}\leq cb_{n}$ holds for all large $n$. If $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ both hold, then we write $a_{n}\asymp b_{n}$. Lastly, let $a_{n}\vee b_{n}=\max\{a_{n},b_{n}\}$.
2 Method
Here, we provide the details for constructing the bootstrap sample $\mathcal{M}_{n}^{*}$, as well as simultaneous confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for the coordinate-wise means $\mu_{1},\dots,\mu_{p}$. Further applications of these intervals to various testing problems will be covered later in Section 4.
In addition to the observations $X_{1},\dots,X_{n}$ discussed above, let $X_{n+1},\dots,X_{n+m_{n}}$ denote an independent set of i.i.d. hold-out observations generated from the same distribution. For simplicity, the number of hold-out observations $m_{n}$ is assumed to be even, and in all of our numerical experiments, we will take $m_{n}$ to be about 10% of $n$. The hold-out observations are used to construct robust estimators $\tilde{X}_{1},\dots,\tilde{X}_{p}$ and $\hat{\sigma}_{1}^{2},\dots,\hat{\sigma}_{p}^{2}$ for the coordinate-wise means and variances, which are the only ingredients for generating $\mathcal{M}_{n}^{*}$ that were not addressed previously in Section 1. Taking an mom approach, we partition the hold-out indices $\{n+1,\dots,n+m_{n}\}$ into $b_{n}$ blocks $\mathcal{B}_{1},\dots,\mathcal{B}_{b_{n}}$, with each block containing an even number of $\ell_{n}$ indices such that $m_{n}=\ell_{n}b_{n}$. More specifically, let $\mathcal{B}_{1}=\{n+1,\dots,n+\ell_{n}\}$, $\mathcal{B}_{2}=\{n+\ell_{n}+1,\dots,n+2\ell_{n}\}$, and so on. For the $l$th block, let
$\bar{X}_{j}(l)=\frac{1}{\ell_{n}}\sum_{i\in\mathcal{B}_{l}}X_{ij}$ (3)
denote the block-wise sample mean of the $j$th coordinate, and define the mom estimator of $\mu_{j}$ as
$\small\tilde{X}_{j}=\textup{median}(\bar{X}_{j}(1),\dots,\bar{X}_{j}(b_{n})).$ (4)
Likewise, we construct an mom estimate for each $\sigma_{j}^{2}$ along similar lines. The $l$th blockwise estimate for $\sigma_{j}^{2}$ is obtained by averaging the squared differences of $\ell_{n}/2$ pairs of observations
$\small\bar{\sigma}_{j}^{2}(l)=\frac{1}{\ell_{n}/2}\sum_{\begin{subarray}{c}i,i^{\prime}\in\mathcal{B}_{l}\\ i^{\prime}-i=\ell_{n}/2\end{subarray}}\frac{1}{2}(X_{ij}-X_{i^{\prime}j})^{2},$ (5)
and then $\hat{\sigma}_{j}^{2}$ is defined to be the median of the block-wise estimates
$\small\hat{\sigma}_{j}^{2}=\textup{median}(\bar{\sigma}_{j}^{2}(1),\dots,\bar{\sigma}_{j}^{2}(b_{n})).$ (6)
Next, we turn to the construction of simultaneous confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\mu_{1},\dots,\mu_{p}$. Let $1-\alpha$ denote the nominal simultaneous coverage probability, and let $\hat{q}_{+}(1-\alpha/2)$ denote the empirical $(1-\alpha/2)$-quantile of a collection of bootstrap samples generated in the manner of $\mathcal{M}_{n}^{*}$. Also, let $\underline{\mathcal{M}}_{n}^{*}$ denote the counterpart of $\mathcal{M}_{n}^{*}$ that is obtained by replacing $\max_{1\leq j\leq p}$ with $\min_{1\leq j\leq p}$ in equation (2), and let $\hat{q}_{-}(\alpha/2)$ denote the empirical $(\alpha/2)$-quantile of a collection of bootstrap samples generated in the manner of $\underline{\mathcal{M}}_{n}^{*}$. In this notation, the confidence interval $\hat{\mathcal{I}}_{j}$ is defined by
$\hat{\mathcal{I}}_{j}=\bigg\{x\in\mathbb{R}\,\mathrel{\mathop{\ordinarycolon}}\,\hat{q}_{-}(\alpha/2)\leq\frac{1}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}\sum_{i=1}^{n}\varphi_{\hat{t}_{j}}(X_{ij}-x)\leq\hat{q}_{+}(1-\alpha/2)\bigg\}.$ (7)
Due to the fact that the functions $\varphi_{\hat{t}_{1}}(\cdot),\dots,\varphi_{\hat{t}_{p}}(\cdot)$ are monotone, it is straightforward to compute all the endpoints of $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$.
To comment on the role of the partial standardization parameter $\tau\in[0,1]$, it provides a way to balance two opposing effects that occur in the extreme cases when $\tau$ is equal to 0 or 1. When $\tau=0$, all of the intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ have the same width, which is clearly undesirable when the covariates fluctuate over different scales. Alternatively, when $\tau=1$, all of the covariates will be on approximately “equal footing”, which will tend to make the max statistic $\mathcal{M}_{n}$ sensitive to all $p$ dimensions. This is undesirable in high-dimensional situations where the covariates fluctuate over different scales, because it eliminates a form of low-dimensional structure that can simplify the behavior of $\mathcal{M}_{n}$ when $\tau<1$. To see this, consider a case where $\tau=0$ and $\sigma_{1},\dots,\sigma_{d}$ are much larger than $\sigma_{d+1},\dots,\sigma_{p}$ for some $d\ll p$. In this case, the maximizing index for $\mathcal{M}_{n}$ is likely to reside in the small subset $\{1,\dots,d\}\subset\{1,\dots,p\}$. Thus, the behavior of $\mathcal{M}_{n}$ will be mainly governed by the first $d$ covariates, which intuitively reduces the effective dimension of the problem of approximating the distribution of $\mathcal{M}_{n}$. Accordingly, as was originally proposed in (lopes2020bootstrapping), it is natural to select an intermediate value of $\tau$ between 0 and 1 that can mitigate the unwanted effects that occur at $\tau\in\{0,1\}$.
3 Theory
Our theoretical analysis is framed in terms of a sequence of models that are implicitly embedded in a triangular array whose rows are indexed by $n$. In this context, all model parameters are allowed to vary with $n$, except when stated otherwise. In particular, the dimension $p=p(n)$ is regarded as a function of $n$, and hence, if a quantity does not depend on $n$, then it does not depend on $p$ either.
To state our model assumptions, recall that the sorted coordinate-wise variances of $X_{1}$ are denoted as $\sigma_{(1)}^{2}\geq\cdots\geq\sigma_{(p)}^{2}$, and for any $d\in\{1,\dots,p\}$, let $J(d)$ be a set of $d$ indices in $\{1,\dots,p\}$ that satisfies $\{\sigma_{j}^{2}\,|\,j\in J(d)\}=\{\sigma_{(1)}^{2},\dots,\sigma_{(d)}^{2}\}$. In addition, let $R(d)$ denote the $d\times d$ correlation matrix associated with the covariates $\{X_{1j}\}_{j\in J(d)}$.
Assumption 1.
The observations $X_{1},\dots,X_{n+m_{n}}\in\mathbb{R}^{p}$ are i.i.d., and there are constants $C\geq 1$, $\beta>0$, and $\delta\geq\epsilon>0$ not depending on $n$ such that the following conditions hold:
  1. (i)
    For all $v\in\mathbb{R}^{p}$, $\big\|\big\langle v,X_{1}-\mathbf{E}(X_{1})\big\rangle\big\|_{L^{4+\delta}}\,\leq\,C\big\|\big\langle v,X_{1}-\mathbf{E}(X_{1})\big\rangle\big\|_{L^{2}}$ holds.
  2. (ii)
    For all $j=1,\dots,p$, the random variable $X_{1j}/\sigma_{j}$ has a Lebesgue density $f_{j}$ such that $\|f_{j}\|_{L^{\infty}}\leq C$.
  3. (iii)
    For all $j=1,\dots,p$, the inequalities $\frac{1}{C}\sigma_{(1)}^{2}j^{-2\beta}\leq\sigma_{(j)}^{2}\leq C\sigma_{(1)}^{2}j^{-2\beta}$ hold.
  4. (iv)
    If $l_{n}=\big\lceil n^{\frac{\epsilon}{24(\beta\vee 1)}}\!\wedge p\big\rceil$, then $\|R(l_{n})\|_{F}^{2}\leq Cl_{n}^{2-\frac{1}{C}}$.
Remarks. All of the conditions in Assumption 1 are invariant to shifting $X_{1}\mapsto X_{1}+v$ for fixed $v\in\mathbb{R}^{p}$, and scaling $X_{1}\mapsto cX_{1}$ for fixed $c\neq 0$. The following paragraphs provide several examples that address each of the conditions ((i))-((iv)). Also, it is straightforward to combine the examples to construct a wide assortment of data-generating distributions that satisfy all of the conditions in Assumption 1 simultaneously.
Examples of $L^{4+\delta}$-$L^{2}$ moment equivalence. In recent years, moment assumptions similar to condition ((i)) have been adopted in many analyses of robust statistical methods for high-dimensional data lugosi2019mean; ke2019user; mendelson2020robust; roy2021empirical; abdalla2022covariance. As shown in Proposition 1 below, condition ((i)) is compatible with the classes of elliptical and separable models (also known as independent component models), which are widely used in areas such as multivariate analysis, random matrix theory, and signal processing (kotz2019continuous; Bai:2010; Comon:2010).
To be precise, we say that an observation $X_{1}$ with mean $\mu$ and covariance matrix $\Sigma$ has an elliptical distribution if it can be represented as $X_{1}=\mu+\eta_{1}\Sigma^{1/2}U_{1}$, where $U_{1}\in\mathbb{R}^{p}$ is uniformly distributed on the unit sphere, and $\eta_{1}$ is a non-negative scalar random variable that is independent of $U_{1}$ and normalized by $\mathbf{E}(\eta_{1}^{2})=p$. On the other hand, we say that $X_{1}$ has a separable distribution if it can be represented as $X_{1}=\mu+\Sigma^{1/2}\zeta_{1}$, where $\zeta_{1}=(\zeta_{11},\dots,\zeta_{1p})$ has i.i.d. entries with $\mathbf{E}(\zeta_{11})=0$, and $\textup{var}(\zeta_{11})=1$.
Proposition 1.
Conditions (i) and (ii) hold simultaneously if one of the following two conditions holds for some $\delta>0$ that is fixed with respect to $n$.
  • (1)
    The observation $X_{1}$ is drawn from an elliptical distribution such that $\|\eta_{1}\|_{L^{4+\delta}}\lesssim\sqrt{p}$, and the random variable $X_{11}/\sigma_{1}$ has a Lebesgue density $f_{1}$ such that $\|f_{1}\|_{L^{\infty}}\lesssim 1$.
  • (2)
    The observation $X_{1}$ is drawn from a separable distribution with
    $\max_{1\leq j\leq p}\|\zeta_{1j}\|_{L^{4+\delta}}$ $\lesssim$ 11, and each random variable $\zeta_{1j}$ has a Lebesgue density $g_{j}$ such that $\max_{1\leq j\leq p}\|g_{j}\|_{L^{\infty}}\lesssim 1$.
The proof is provided in Appendix G.
Examples of variance decay. There are a variety of settings where the sorted coordinate-wise variances $\sigma_{(1)}^{2}\geq\cdots\geq\sigma_{(p)}^{2}$ naturally exhibit a decay profile.
Principal components analysis. In the context of principal components analysis, it is common to assume that the sorted eigenvalues $\lambda_{1}(\Sigma)\geq\cdots\geq\lambda_{p}(\Sigma)$ of $\Sigma$ satisfy $\lambda_{j}(\Sigma)\lesssim j^{-\gamma}$ for some $\gamma>0$, and it can be shown that this implies $\sigma_{(j)}^{2}\lesssim j^{-2\beta}$ for some other decay parameter $\beta>0$ (lopes2020bootstrapping, Proposition 2.1).
Mean-variance proportionality. Another scenario where variance decay arises is when the coordinate-wise means and variances are connected by a proportionality relationship of the form $\sigma_{j}^{2}\propto|\mu_{j}|^{\gamma}$, for some fixed exponent $\gamma>0$. This occurs within many sub-families of distributions, including Gamma, Weibull, inverse Gaussian, and Pareto. In applications that involve sparse modelling of high-dimensional mean vectors, a classical assumption is that the sorted coordinate-wise means have a decay profile (Johnstone), and thus, when such a proportionality relationship holds, it follows that variance decay must also occur.
Functional data analysis. One more set of examples is related to functional data analysis, where function-valued observations $\Psi_{1},\dots,\Psi_{n}$ in a Hilbert space are often studied through their projections under a finite number of orthonormal basis functions $\{\phi_{j}\}_{1\leq j\leq p}$. That is, the $i$th projected observation has the form $X_{i}=(\langle\Psi_{i},\phi_{1}\rangle,\dots,\langle\Psi_{i},\phi_{p}\rangle)\in\mathbb{R}^{p}$. In connection with our work, the important point is that under standard assumptions in functional data analysis, it can be shown that the sorted coordinate-wise variances of $X_{i}$ have a decay profile (lopes2020bootstrapping). In fact, this occurs even when the random functions $\Psi_{1},\dots,\Psi_{n}$ have rough sample paths, which we illustrate empirically in Figure 2.
Examples of correlation matrices. To interpret the condition ((iv)), it should be noted that the inequality $\|R(l_{n})\|_{F}^{2}\leq l_{n}^{2}$ always holds, since $\|A\|_{F}^{2}\leq\textup{tr}(A)^{2}$ holds for any positive semidefinite matrix $A$. So, in this sense, condition ((iv)) is quite mild, as $C$ may be taken to be arbitrarily large. Moreover, the correlation structure of the variables indexed by $\{1,\dots,p\}\setminus J(l_{n})$ is completely unrestricted. With regard to the constant 24 appearing in the definition of $l_{n}$, it has no special importance, and is used for theoretical convenience. Below, we describe several classes of $p\times p$ correlation matrices $R=R(p)$ for which the sub-matrix $R(l_{n})$ satisfies condition ((iv)).
Decaying correlation functions. Let $\rho\mathrel{\mathop{\ordinarycolon}}[0,\infty)\rightarrow[0,1]$ be any continuous convex function satisfying $\rho(0)=1$, and $\rho(t)\leq ct^{-\gamma}$ for some fixed constants $c>0,\gamma>0$, and all $t\geq 0$. By Pólya’s criterion (polya1949remarks), a matrix whose $ij$ entry is defined by $\rho(|i-j|)$ is a correlation matrix that satisfies condition ((iv)). Correlation matrices of this type include many well-known examples, such as those of the autoregressive and banded types, e.g.,  $R_{ij}=r^{|i-j|}$ for some fixed $r\in(0,1)$, and $R_{ij}=\max\big\{0,1-\frac{|i-j|}{b}\big\}$ for some fixed $b>0$.
Diverging operator norm. If the operator norm of $R$ satisfies $\|R\|_{\textup{op}}\leq Cl_{n}^{\frac{1}{2}-\frac{1}{2C}}$, then Assumption 1((iv)) holds. This can be seen by noting that $\|R(l_{n})\|_{F}^{2}\leq l_{n}\|R(l_{n})\|_{\textup{op}}^{2}\leq l_{n}\|R\|_{\textup{op}}^{2}$. In particular, since $l_{n}$ increases with $n$ and $p$, this shows that condition ((iv)) can hold even when the operator norm of $R$ diverges asymptotically.
Block structure. Suppose that $R$ is formed by concatenating $k$ correlation matrices along its diagonal, with sizes $\nu_{1}\times\nu_{1},\dots,\nu_{k}\times\nu_{k}$, so that $\nu_{1}+\cdots+\nu_{k}=p$. If the condition $\max\{\nu_{1},\dots,\nu_{k}\}\leq Cl_{n}^{1-\frac{1}{C}}$ holds, then so does condition ((iv)). This follows from the observation that no row of $R$ can have a squared $\ell_{2}$ norm larger than $\max\{\nu_{1},\dots,\nu_{k}\}$, and so $\|R(l_{n})\|_{F}^{2}\leq l_{n}\max\{\nu_{1},\dots,\nu_{k}\}$.
Convex combinations and permutations. If $R$ and $R^{\prime}$ denote any correlation matrices corresponding to the previous examples, then for any $t\in[0,1]$, the correlation matrix $tR+(1-t)R^{\prime}$ satisfies condition ((iv)). Furthermore, if $\Pi$ is a $p\times p$ permutation matrix, then $\Pi R\Pi^{\top}$ is also a correlation matrix that satisfies condition ((iv)). These operations considerably extend the examples that have decaying correlation functions or block structure.
The following theorem is our main result.
Theorem 1.
Fix any constant $\tau\in[0,1)$ with respect to $n$, and suppose that Assumption 1 holds with the values of $\delta\geq\epsilon>0$ stated there. In addition, suppose that the hold-out set consists of $m_{n}\asymp n$ observations that are partitioned into $b_{n}\asymp\log(n)$ blocks. Then, there is a constant $c>0$ not depending on $n$ such that the event
$\sup_{s\in\mathbb{R}}\Big|\mathbf{P}(\mathcal{M}_{n}\leq s)-\mathbf{P}(\mathcal{M}_{n}^{*}\leq s|X)\Big|\ \leq\ c\ \!n^{-\frac{1}{2}+\epsilon}$
occurs with probability at least $1-c\ \!n^{-\delta/4}$.
[…]
The proof also analyzes several effects that arise from heavy-tailed covariates and the structure of the robust max statistic $\mathcal{M}_{n}$. A particularly important example of such an effect is the bias that is introduced by the truncation functions $\varphi_{\hat{t}_{j}}$, and this is addressed in the proofs of Propositions 4 and 7, as well as Lemmas  3,  4, and 7. Furthermore, our work accounts for the fluctuations of the robust mom estimates $\tilde{X}_{j}$ and $\hat{\sigma}_{j}$ that are used to partially standardize $\mathcal{M}_{n}^{*}$, and this is done in Lemmas 12-15.
4 Numerical experiments
This section addresses the practical performance of the proposed method in the contexts of Euclidean and functional data with heavy tails. In addition, we include the performance comparisons with the method for robust inference proposed in (Fan:2023).
4.1 Euclidean data
Here, we consider the task of constructing simultaneous confidence intervals for the entries of the mean vector $(\mu_{1},\dots,\mu_{p})=\mathbf{E}(X_{1})$, based on i.i.d. observations $X_{1},\dots,X_{n+m_{n}}\in\mathbb{R}^{p}$.
Design of experiments. For each pair $(n+m_{n},p)$ in the set $\{500\}\times\{100,500,1000\}$, and each of the data-generating distributions described below, we performed 500 Monte Carlo trials. Due to the fact that the proposed method and the method in (Fan:2023) are shift invariant, the mean vector $(\mu_{1},\dots,\mu_{p})$ was always chosen to be the zero vector without loss of generality. In all trials, the proposed confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ were constructed according to (7), using 500 bootstrap samples, and using choices of 90% and 95% for the nominal simultaneous coverage probability $1-\alpha$. Also, the proposed method was always applied with truncation parameters set to $\hat{t}_{j}=\sqrt{n}\hat{\sigma}_{j}$ for all $j=1,\dots,p$, and with the partial standardization parameter set to $\tau=0.9$. Lastly, in all trials, the number of hold-out observations was set to $m_{n}=50$, and the block length for the mom estimates was set to $\ell_{n}=10$.
Data-generating distributions. The data were generated in four ways, based on an elliptical distribution and a separable distribution with two choices of covariance matrices.
Elliptical distribution. The elliptical observations have the form $X_{1}=\mu+\eta_{1}\Sigma^{1/2}U_{1}$, where $U_{1}\in\mathbb{R}^{p}$ is uniformly distributed on the unit sphere, and $\eta_{1}\geq 0$ is a random variable that is independent of $U_{1}$ such that $3\eta_{1}^{2}/(2p)$ follows an $F$ distribution with $p$ and 6 degrees of freedom. This distribution for $X_{1}$ is more commonly known as a multivariate $t$-distribution on 6 degrees of freedom (muirhead2009aspects).
Separable distribution. The separable observations have the form $X_{1}=\mu+\Sigma^{1/2}\zeta_{1}$, where $\zeta_{1}=(\zeta_{11},\dots,\zeta_{1p})$ has i.i.d. entries that are standardized Pareto random variables. Specifically, $\zeta_{11}=(\omega_{11}-\mathbf{E}(\omega_{11}))/\sqrt{\textup{var}(\omega_{11})}$, where $\omega_{11}$ is drawn from a Pareto distribution whose density is given by $x\mapsto 6x^{-7}1\{x\geq 1\}$.
Covariance matrices. For both the elliptical and separable distributions, we constructed the covariance matrix of $X_{1}$ in the form $\Sigma=D^{1/2}RD^{1/2}$, where $D=\textup{diag}(\textup{var}(X_{11}),\dots,\textup{var}(X_{1p}))$, and $R$ is the correlation matrix of $X_{1}$. The correlation matrix was chosen to be one of the following two types
$\small R_{ij}=\begin{cases}0.5^{|i-j|}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{(autoregressive)}\\ \textstyle 1\{i=j\}+\frac{1\{i\neq j\}}{4(i-j)^{2}}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{(algebraic decay).}\end{cases}$
In all cases, the matrix $D$ was chosen to satisfy $D_{jj}^{1/2}=j^{-1/2}$ for all $j=1,\dots,p$, ensuring that the entries of $X_{1}$ have variance decay.
Discussion of results. In Table 1, we report empirical simultaneous coverage probabilities and width measures, for both the proposed method (denoted PM) and the method based on the Hodges-Lehmann estimator (denoted HL) developed in (Fan:2023). The simultaneous coverage probabilities were computed as the fraction of the 500 Monte Carlo trials in which all $p$ intervals of a given method contained the corresponding parameters $\mu_{1},\dots,\mu_{p}$. The width measure was computed as the median width of the $p$ intervals, averaged over the 500 trials, and it is shown in parentheses below the simultaneous coverage probabilities.
Across all of the settings, PM delivers simultaneous coverage probabilities that are within about 1 or 2 percent of the nominal level, demonstrating that it is reliably calibrated. On the other hand, HL is only well calibrated in the cases of elliptical models, where the covariates have distributions that are symmetric around 0. In the cases of separable models where the covariates have distributions that are not symmetric around 0, the simultaneous coverage probabilities of HL are far from the nominal level. This can be explained by the fact that HL is designed for simultaneous inference on the coordinate-wise pseudomedians—which may differ from the coordinate-wise means if the covariates have asymmetric distributions. By contrast, PM has a scope of application for inference on coordinate-wise means that is not limited by asymmetric distributions.
With regard to the width measure, Table 1 shows that the intervals produced by PM are much tighter than those produced by HL, sometimes by a factor of 4 or more. This is to be expected, because the HL intervals have equal widths across all $p$ coordinates, whereas PM adapts to the scale of each covariate and takes advantage of variance decay. This difference between the methods is also reflected in another pattern—which is that the width measure for PM decreases as $p$ increases, whereas the width measure for HL stays essentially constant as $p$ increases, since HL uses unstandardized covariates.
Table 1: Comparison of simultaneous coverage probability and confidence interval width: PM refers to the proposed method, and HL refers to the method based on the Hodges-Lehmann estimator proposed in (Fan:2023).
$R$ Distribution $\alpha$ $p=100$ $p=500$ $p=1000$
PM HL PM HL PM HL
auto- regressive elliptical 0.05 0.942 0.942 0.956 0.938 0.962 0.962
(0.057) (0.165) (0.031) (0.165) (0.023) (0.166)
0.1 0.914 0.892 0.918 0.900 0.908 0.916
(0.053) (0.142) (0.029) (0.142) (0.022) (0.142)
separable 0.05 0.956 0.006 0.946 0.002 0.954 0
(0.065) (0.143) (0.035) (0.144) (0.027) (0.143)
0.1 0.922 0 0.904 0.002 0.910 0
(0.059) (0.123) (0.033) (0.124) (0.025) (0.123)
algebraic decay elliptical 0.05 0.946 0.938 0.942 0.942 0.958 0.962
(0.058) (0.167) (0.031) (0.166) (0.023) (0.166)
0.1 0.914 0.890 0.910 0.898 0.916 0.904
(0.053) (0.144) (0.029) (0.143) (0.022) (0.144)
separable 0.05 0.968 0 0.954 0 0.958 0
(0.067) (0.141) (0.037) (0.142) (0.028) (0.141)
0.1 0.922 0 0.910 0 0.912 0
(0.062) (0.122) (0.034) (0.123) (0.026) (0.122)
4.2 Functional data
Beyond heavy-tailed Euclidean data, our approach to inference with robust max statistics can be applied to heavy-tailed functional data. Below, we study the problem of detecting a non-zero drift in functional observations that arise from geometric Brownian motion (GBM). This is a heavy-tailed stochastic process in the sense that its marginals are lognormal, which is a standard example of heavy-tailed univariate distribution (Foss; nair). The sample paths of GBM present additional challenges from the standpoint of functional data analysis, because they are rough. Furthermore, GBM is of broad interest in financial applications, where it is widely used for modelling securities prices (Oksendal).
Problem formulation. A sample path of GBM on the unit interval has the form $t\mapsto\exp\big((\mu_{0}-\varsigma_{0}^{2}/2)t+\varsigma_{0}W(t)\big)$ for $t\in[0,1]$, where $W(t)$ is a standard Brownian motion, $\mu_{0}\in\mathbb{R}$ is the drift parameter, and $\varsigma_{0}^{2}\geq 0$ is the volatility parameter. To formalize the detection of non-zero drift in a way that allows us to incorporate natural alternative hypotheses, we will allow for more general sample paths $S(t)$ of the form
$\small S(t)=\exp\big((h\mu(t)-\varsigma_{0}^{2}/2)t+\varsigma_{0}W(t)\big),$ (8)
where $\mu(t)$ is a fixed real-valued function on $[0,1]$ such that $S(t)$ resides in $L^{2}[0,1]$ almost surely, and $h\geq 0$ is a fixed parameter that measures the “distance” from the null hypothesis of zero drift that occurs when $h=0$. The parameters $\mu(t)$ and $\varsigma_{0}$ are treated as unknown. Under these conditions, we are interested in using a dataset $S_{1}(t),\dots,S_{n+m_{n}}(t)$ of i.i.d. samples of $S(t)$ to address the hypothesis testing problem
$\small\mathsf{H}_{0}\mathrel{\mathop{\ordinarycolon}}h=0\text{ \ \ \ vs. \ \ \ }\mathsf{H}_{1}\mathrel{\mathop{\ordinarycolon}}h>0.$ (9)
In particular, different choices of the function $\mu(t)$ correspond to different alternatives, and later on, we will present numerical results for several choices.
Testing procedure. The sample path formula (8) implies that $\mathbf{E}(S(t))=\exp(h\mu(t)t)$ for all $t\in[0,1]$. For this reason, our procedure will seek to detect whether or not the function $\mathbf{E}(S(t))-1$ is identically 0. This will be done by expanding $\mathbf{E}(S(t))-1$ in the form $\sum_{j=1}^{\infty}\beta_{j}\phi_{j}(t)$, where $\{\phi_{j}(t)\}_{j\geq 1}$ is the Fourier cosine basis for $L^{2}[0,1]$. To proceed, note that $\mathbf{E}(S(t))-1$ is equal to the zero function in the $L^{2}[0,1]$ sense if and only if $\beta_{j}=0$ for all $j\geq 1$. This motivates a procedure based on testing the simultaneous hypotheses
$\small\mathsf{H}_{0,j}\mathrel{\mathop{\ordinarycolon}}\beta_{j}=0\text{ \ \ \ for \ \ \ }j=1,\dots,p,$ (10)
where $p$ is an integer large enough so that the coefficients $\beta_{p+1},\beta_{p+2},\dots,$ are negligible for practical purposes. In particular, $\mathsf{H}_{0}$ implies that $\mathsf{H}_{0,1},\dots,\mathsf{H}_{0,p}$ hold simultaneously, and so a procedure that controls the simultaneous type I error rate for these hypotheses leads to one that controls the ordinary type I error rate for $\mathsf{H}_{0}$.
If we define $X_{i}\in\mathbb{R}^{p}$ to contain the first $p$ coefficients of $S_{i}(t)-1$ with respect to $\{\phi_{j}(t)\}_{j\geq 1}$, then it follows that $\mathbf{E}(X_{i})=(\beta_{1},\dots,\beta_{p})$ for every $i=1,\dots,n+m_{n}$. Hence, we may apply our proposed method from Section 2 to $X_{1},\dots,X_{n+m_{n}}$ in order to construct confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\beta_{1},\dots,\beta_{p}$ with a nominal simultaneous coverage probability of $1-\alpha$. Altogether, this means that if we reject $\mathsf{H}_{0}$ when any of these intervals exclude 0, then this rejection rule corresponds to a test with a nominal level of at most $\alpha$.
Design of experiments. To construct four natural choices of the pair $(\mu(t),\varsigma_{0})$, we used historical data for the stocks of Apple, Nvidia, Moderna, and JPMorgan, to be described later in the paragraph labeled ‘preparation of stock data’. With regard to the choice of $\varsigma_{0}^{2}$, note that if a stock price is modeled as a realization of $S(t)$, then the pointwise variance of the cumulative log return is $\textup{var}((h\mu(t)-\varsigma_{0}^{2}/2)t+\varsigma_{0}W(t))=\varsigma_{0}^{2}t$, and the time average of this quantity over the unit interval is $\varsigma_{0}^{2}/2$. (See (ruppertstatistics) for additional background.) Accordingly, for each of the four stocks, we selected $\varsigma_{0}$ such that $\varsigma_{0}^{2}/2=\int_{0}^{1}s^{2}(t)\,\mathrm{d}t$, where $s^{2}(t)$ is the sample pointwise variance of the cumulative log return of the stock over many disjoint periods of unit length. Next, the selection of $\mu(t)$ was motivated by the fact that if a stock price is modeled with $S(t)$, and if $h=1$, then the pointwise expected cumulative log return is $(\mu(t)-\varsigma_{0}^{2}/2)t$. For a given value of $\varsigma_{0}$, we defined $\mu(t)$ to satisfy $(\mu(t)-\varsigma_{0}^{2}/2)t=\bar{R}(t)$, where $\bar{R}(t)$ is the sample average of the cumulative log return curves of a given stock over many disjoint periods of unit length.
For each of the four choices of $(\mu(t),\varsigma_{0})$, and each value of $h$ in an equispaced grid, the following procedure was repeated in 500 Monte Carlo trials. We generated i.i.d. realizations $S_{1}(t),\dots,S_{n+m_{n}}(t)$ of the sample path defined in (8) with $n+m_{n}=300$ and $m_{n}=30$. For each of these sets of functional observations, we constructed simultaneous $(1-\alpha)$-confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for the parameters $\beta_{1},\dots,\beta_{p}$, as described above, with $p=100$ and $\alpha=5\%$. Whenever any of these intervals excluded 0, a rejection was recorded, and the rejection rate among the 500 trials was plotted as a function of $h$ in Figures 3(a)-3(d). The corresponding rejection rate based on the simultaneous confidence intervals proposed in the paper (Fan:2023) was also plotted in the same way. In all four figures, the nominal level of $\alpha=5\%$ is marked with a dashed horizontal line.
To illustrate the characteristics of the simulated functional data, ten realizations of $S(t)$ based on $h=0$ with $\varsigma_{0}$ corresponding to Apple stock are plotted in Figure 2. In the same setting, Figure 2 displays estimates of the sorted values $\sigma_{(1)}\geq\cdots\geq\sigma_{(p)}$, where $\sigma_{j}^{2}$ is the variance of the $j$th Fourier coefficient of $S_{1}(t)$. In particular, Figure 2 shows a clear variance decay profile.
[…]
Figure 1: Representative sample paths of $S(t)=\exp\big((h\mu(t)-\varsigma_{0}^{2}/2)t+\varsigma_{0}W(t)\big)$ when $h=0$, and $\varsigma_{0}$ is selected based on historical price data for Apple stock.
[…]
Figure 2: Estimates of the ordered values $\sigma_{(1)}\geq\cdots\geq\sigma_{(p)}$, where $\sigma_{j}^{2}$ denotes the variance of the $j$th Fourier coefficient of sample paths generated as in Figure 2.
Preparation of stock data. Price data for 4 stocks (Apple, Nvidia, Moderna, and JP Morgan) were collected from the Alpha Vantage database (alphavantage2024) during every minute of trading between March 1, 2024 and March 22, 2024, including pre-market and after-hours trading. The data were divided into 100-minute intervals (normalized to unit length), and each interval was divided into time points $t_{0},t_{1},\ldots,t_{100}$ spaced one minute apart. For a given stock, letting $P(t_{j})$ denote its price at time $t_{j}$, we computed the cumulative log return curve within the interval as $R(t_{j})=\sum_{\ell=1}^{j}\log\big(P(t_{\ell})/P(t_{\ell-1})\big)=\log\big(P(t_{j})/P(t_{0})\big)$, $j=1,\ldots,100$. In this way, we obtained one discretely observed realization of the function $R(t)$ over each 100-minute interval. To promote independence among these functional observations, we only retained them from every other 100-minute interval, ensuring that they are separated by gaps of 100 minutes. We also excluded functional observations that were obtained from intervals with missing data.
[…]
(a) Apple stock
[…]
(b) Nvidia stock
[…]
(c) JP Morgan stock
[…]
(d) Moderna stock
Figure 3: The panels compare the rejection rates of the methods PM and HL when functional observations are generated in the form (8) and the parameters $(\mu(t),\varsigma_{0})$ are based on historical stock price data for Apple, Nvidia, JP Morgan, and Moderna.
Discussion of results. In Figures 3(a)-3(d), the rejection rate curves for the proposed method are labeled by PM, and the corresponding curves based on the intervals proposed in (Fan:2023) are labeled HL. Recall that $h=0$ under the null hypothesis $\mathsf{H}_{0}$, and so the value of a curve at $h=0$ represents the empirical level. It is clear that in all four panels of Figure 3, the empirical levels of both methods closely match the nominal level of $\alpha=5\%$, marked with a dashed horizontal line.
However, the methods differ markedly in terms of power, which is represented by the values of the curves at $h>0$. In the three settings based on the stock data of Apple, Nvidia, and JP Morgan, the power of PM increases steadily with $h$, whereas the power of HL stays relatively flat. In the setting based on Moderna stock data, the two methods are more competitive, but even here, the power of PM is still noticeably higher for most values of $h$. The power advantage of PM is understandable in light of the characteristics of the simultaneous confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\beta_{1},\dots,\beta_{p}$. (Recall that both methods reject the null hypothesis $\mathsf{H}_{0}\mathrel{\mathop{\ordinarycolon}}h=0$ whenever any of their associated intervals exclude 0.) As was observed in Section 4.1, the intervals produced by PM tend to be tighter than those produced by HL, and tighter intervals make it easier to exclude 0, resulting in higher power.
Acknowledgement
We are grateful to Mengxin Yu for generously providing the code for the HL method.
Supplementary Material
Robust Max Statistics for High-Dimensional Inference
Appendix A introduces preliminary material not covered in the main text. Appendix B outlines the proof of Theorem 1, and the main supporting arguments are given in Appendices C-E. Appendix F contains technical results on median-of-means estimators. Appendix G proves Proposition 1. Lastly, Appendix H contains various background results.
Appendix A Preliminaries for supplementary material
Notation. The distribution of any random variable $U$ is denoted as $\mathcal{L}(U)$. We write $\mathcal{L}(U|X)$ to refer to the conditional distribution of $U$ given both the hold-out and non-hold-out sets of observations, whereas we write $\mathcal{L}(U|X^{\prime})$ to refer to the conditional distribution of $U$ given the hold-out set only. Similarly, we use $\mathbf{P}(\cdot|X)$ and $\mathbf{P}(\cdot|X^{\prime})$ to denote conditional probabilities in the two cases just mentioned, and we use $\|\cdot\|_{L^{q}|X}$ and $\|\cdot\|_{L^{q}|X^{\prime}}$ to denote the corresponding conditional $L^{q}$ norms.
For any $d\in\{1,\ldots,p\}$, recall that $J(d)$ denotes a set of indices corresponding to the $d$ largest values among $\{\sigma_{1},\ldots,\sigma_{p}\}$. That is, $\{\sigma_{(1)},\ldots,\sigma_{(d)}\}=\{\sigma_{j}|j\in J(d)\}$. Letting $l_{n}$ be as in Assumption 1, define the integer
$k_{n}=l_{n}^{5}\wedge p,$
which always satisfies $1\leq l_{n}\leq k_{n}\leq p$. For each $d\in\{1,\dots,p\}$, let
$M_{d}(X)=\max_{j\in J(d)}\frac{1}{\sigma_{j}^{\tau}\sqrt{n}}\sum_{i=1}^{n}(X_{ij}-\mu_{j}).$ (11)
Letting $G(X)=(G_{1}(X),\ldots,G_{p}(X))$ be a centered Gaussian random vector with the same covariance matrix as $X_{1}$, the Gaussian counterpart of $M_{d}(X)$ is defined as
$\tilde{M}_{d}(X)=\max_{j\in J(d)}\frac{G_{j}(X)}{\sigma_{j}^{\tau}}.$
Next, for each $i\in\{1,\dots,n\}$ and $j,d\in\{1,\dots,p\}$, define
[…]
as well as
$\hat{Y}_{ij}=\varphi_{\hat{t}_{j}}(X_{ij}-\mu_{j})\text{ \ \ \ and \ \ \ }M_{d}(\hat{Y})=\max_{j\in J(d)}\frac{1}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\sum_{i=1}^{n}\hat{Y}_{ij}.$
Let $\xi_{1},\dots,\xi_{n}$ be i.i.d. standard Gaussian random variables, generated independently of $X_{1},\ldots,X_{n+m_{n}}$, and define
$M_{d}^{*}(X)=\max_{j\in J(d)}\frac{1}{\sigma_{j}^{\tau}\sqrt{n}}\sum_{i=1}^{n}\xi_{i}(X_{ij}-\bar{X}_{j}).$ (12)
Let $\tilde{X}_{j}$ denote the median-of-means estimator for $\mu_{j}$ described in Section 2 and define
$\hat{Z}_{ij}$ $=\varphi_{\hat{t}_{j}}(X_{ij}-\tilde{X}_{j})\text{ \ \ \ \ and \ \ \ }M_{d}^{*}(\hat{Z})=\max_{j\in J(d)}\frac{1}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j}).$
where we let $\hat{Z}_{j}=\frac{1}{n}\sum_{i=1}^{n}\hat{Z}_{ij}$.
Frequently-used inequalities. As a shorthand for the Kolmogorov metric between generic random variables $U$ and $V$ we write
$d_{\mathrm{K}}(\mathcal{L}(U),\mathcal{L}(V))=\sup_{t\in\mathbb{R}}|\mathbf{P}(U\leq t)-\mathbf{P}(V\leq t)|.$
We will often use the following two basic inequalities that hold for any random variables $U$ and $V$, and any number $s>0$,
$d_{\mathrm{K}}(\mathcal{L}(U),\mathcal{L}(V))\ \leq\ \sup_{t\in\mathbb{R}}\mathbf{P}(|V-t|\leq s)\,+\,\mathbf{P}(|U-V|>s),$ (13)
and
$\sup_{t\in\mathbb{R}}\mathbf{P}(|U-t|\leq s)\ \leq\ \sup_{t\in\mathbb{R}}\mathbf{P}(|V-t|\leq s)+2d_{\textup{K}}(\mathcal{L}(U),\mathcal{L}(V)).$ (14)
If $g$ is a centered Gaussian random variable and $q\geq 1$, then there is an absolute constant $c>0$ such that
$\|g\|_{L^{q}}\ \leq\ c\sqrt{q}\|g\|_{L^{2}},$ (15)
as recorded in [vershynin2012introduction, Eqn. 2.11]. When referring to Chebyshev’s inequality, we will typically use it in the following form for a generic random variable $U$,
$\mathbf{P}(|U|\geq e\|U\|_{L^{q}})\leq e^{-q}.$ (16)
For any random variables $U_{1},\dots,U_{p}$, we have
$\Big\|\max_{1\leq j\leq p}U_{j}\Big\|_{L^{q}}\,\leq\,p^{1/q}\max_{1\leq j\leq p}\|U_{j}\|_{L^{q}}.$ (17)
Lastly, for any two real vectors $(a_{1},\ldots,a_{p})$ and $(b_{1},\ldots,b_{p})$, we have
$\Big|\max_{1\leq j\leq p}a_{j}-\max_{1\leq j\leq p}b_{j}\Big|\leq\max_{1\leq j\leq p}|a_{j}-b_{j}|.$ (18)
Appendix B Proof of Theorem 1
Observe that the left side of the bound in Theorem 1 is given by
$\sup_{s\in\mathbb{R}}\Big|\mathbf{P}(\mathcal{M}_{n}\leq s)-\mathbf{P}(\mathcal{M}_{n}^{*}\leq s|X)\Big|=d_{\textup{K}}\big(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(M_{p}^{*}(\hat{Z})|X)\big).$ (19)
We will bound this distance in three main parts
$d_{\textup{K}}\big(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(M_{p}^{*}(\hat{Z})|X)\big)$ $\ \ \leq\ \ d_{\textup{K}}\big(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(\tilde{M}_{k_{n}}(X))\big)$ (20)
$\ \ \ \ \ \ +\ d_{\textup{K}}\Big(\mathcal{L}(\tilde{M}_{k_{n}}(X)),\mathcal{L}(M_{k_{n}}^{*}(X)|X)\big)$ (21)
$\ \ \ \ \ \ +\ d_{\textup{K}}\big(\mathcal{L}(M_{k_{n}}^{*}(X)|X),\mathcal{L}(M_{p}^{*}(\hat{Z})|X)\big).$ (22)
The three terms on the right side respectively correspond to a Gaussian approximation, a Gaussian comparison, and a bootstrap approximation. These terms are respectively addressed in Proposition 2 of Appendix C, Proposition 6 of Appendix D and Proposition 7 of Appendix E, which show that all the terms are $\mathcal{O}(n^{-\frac{1}{2}+\epsilon})$ with probability at least $1-\mathcal{O}(n^{-\delta/4})$. ∎
Conventions. In the appendices supporting the proof of Theorem 1, we may assume without loss of generality that $\epsilon$ satisfies $\epsilon<1/2$ and that $n\geq c$ for any fixed constant $c>0$, for otherwise the result is trivially true. Also, we will often use $c$ to denote a generic positive constant not depending on $n$, whose value may differ at each appearance. Lastly, we may assume without loss of generality that $(\mu_{1},\dots,\mu_{p})=\mathbf{E}(X_{1})=0$, because the conditions in Assumption 1 are shift invariant, and that $\max_{1\leq j\leq p}\sigma_{j}^{2}=1$, because the Kolmogorov metric is scale invariant. To avoid repetitiveness, these conventions will not be stated explicitly in most of the results presented in the appendices.
Appendix C Gaussian approximation
Proposition 2.
If the conditions of Theorem 1 hold, then
$d_{\textup{K}}(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(\tilde{M}_{k_{n}}(X)))\ \lesssim n^{-\frac{1}{2}+\epsilon}.$
Proof. The proof is based on the decomposition
$d_{\textup{K}}(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(\tilde{M}_{k_{n}}(X)))\leq\mathrm{I}_{n}+\mathrm{II}_{n}+\mathrm{III}_{n},$
where we define
[…]
$\mathrm{II}_{n}=d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(\hat{Y})),\mathcal{L}(M_{k_{n}}(Y))),$
$\mathrm{III}_{n}=d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(Y)),\mathcal{L}(\tilde{M}_{k_{n}}(X))).$
Below, the terms $\mathrm{I}_{n}$, $\mathrm{II}_{n}$, and $\mathrm{III}_{n}$ are shown to be at most of order $n^{-\frac{1}{2}+\epsilon}$ in Propositions 3, 4, and 5 respectively.
Proposition 3.
If the conditions of Theorem 1 hold, then
$d_{\textup{K}}(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(M_{k_{n}}(\hat{Y})))\ \lesssim n^{-\frac{1}{2}+\epsilon}.$
Proof. For any $t\in\mathbb{R}$, define the events
$A(t)=\Big\{\max_{j\in J(k_{n})}\frac{\sum_{i=1}^{n}\hat{Y}_{ij}}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}\leq t\Big\}\quad\text{ and }\quad B(t)=\Big\{\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\hat{Y}_{ij}}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}>t\Big\}.$
It is straightforward to check that for any $t\in\mathbb{R}$, we have
$\Big|\mathbf{P}(M_{p}(\hat{Y})\leq t)-\mathbf{P}(M_{k_{n}}(\hat{Y})\leq t)\Big|\ =\ \mathbf{P}(A(t)\cap B(t)).$
Next, it can be checked that for any real numbers $s_{1,n}$ and $s_{2,n}$ satisfying $s_{1,n}\leq s_{2,n}$, the inclusion
$A(t)\cap B(t)\ \subset\ A(s_{2,n})\cup B(s_{1,n})$
holds simultaneously for all $t\in\mathbb{R}$. Therefore, after taking the supremum over $t\in\mathbb{R}$, we have
$d_{\textup{K}}(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(M_{k_{n}}(\hat{Y})))\ \leq\ \mathbf{P}(A(s_{2,n}))+\mathbf{P}(B(s_{1,n})).$ (23)
Let
$\omega=\frac{\epsilon}{24(\beta\vee 1)C},$ (24)
$d_{n}=\big\lfloor\textstyle\frac{\omega^{2}}{4}{\tt{r}}(R(l_{n}))\vee 2\big\rfloor,$ (25)
where ${\tt{r}}(R(l_{n}))\mathrel{\mathop{\ordinarycolon}}=\textup{tr}(R(l_{n}))^{2}/\|R(l_{n}))\|_{F}^{2}=\frac{l_{n}^{2}}{\|R(l_{n})\|_{F}^{2}}$ is the stable rank of $R(l_{n})$. We will choose $s_{1,n}$ and $s_{2,n}$ according to
$s_{1,n}$ $=c_{1}k_{n}^{-\beta(1-\tau)/2}(\log(n)\vee 2),$ (26)
$s_{2,n}$ $=c_{2}l_{n}^{-\beta(1-\tau)}\sqrt{\log(d_{n})},$
for some constants $c_{1},c_{2}>0$ not depending on $n$. It can also be checked that for any fixed choices of $c_{1}$ and $c_{2}$, the inequality $s_{1,n}\leq s_{2,n}$ holds for all large $n$ due to the definitions of $k_{n},l_{n}$ and $d_{n}$.
To bound $\mathbf{P}(A(s_{2,n}))$, we have
$\mathbf{P}\big(A(s_{2,n})\big)\ \leq\$$\mathbf{P}(\tilde{M}_{k_{n}}(X)\leq s_{2,n})+d_{\mathrm{K}}(\mathcal{L}(M_{k_{n}}(\hat{Y})),\mathcal{L}(\tilde{M}_{k_{n}}(X))),$
where the first term on the right hand side is of order $n^{-1/2}$ by Lemma 1, and the second term is of order $n^{-\frac{1}{2}+\epsilon}$ by Propositions 4 and 5. Lastly, Lemma 2 shows that $\mathbf{P}(B(s_{1,n}))$ is of order $\frac{1}{n}$, which completes the proof.∎
Lemma 1.
Suppose conditions of Theorem 1 hold. Then, there is a constant $c_{2}>0$, not depending on $n$, such that the following bound holds when $s_{2,n}=c_{2}l_{n}^{-\beta(1-\tau)}\sqrt{\log(d_{n})}$,
$\mathbf{P}\Big(\tilde{M}_{k_{n}}(X)\leq s_{2,n}\Big)\lesssim n^{-1/2}.$
Proof. Observe that for any $t\in\mathbb{R}$,
$\mathbf{P}(\tilde{M}_{k_{n}}(X)\leq t)\leq\mathbf{P}(\tilde{M}_{l_{n}}(X)\leq t).$ (27)
Let $(a_{j})_{j\in J(l_{n})}$ and $b$ be positive numbers with $\max_{j\in J(l_{n})}a_{j}\leq b$. For any sequence of random variables $(U_{j})_{j\in J(l_{n})}$ and $t\geq 0$, it is straightforward to check that
$\mathbf{P}\Big(\max_{j\in J(l_{n})}U_{j}\leq t\Big)\leq\mathbf{P}\Big(\max_{j\in J(l_{n})}a_{j}U_{j}\leq bt\Big).$
Consider the choice $a_{j}=\sigma_{j}^{-(1-\tau)}$ and note that under Assumption 1((iii)), there is a constant $c_{0}>0$ not depending on $n$ such that the bound $a_{j}\leq c_{0}l_{n}^{(1-\tau)\beta}\!=\mathrel{\mathop{\ordinarycolon}}b$ holds for all $j\in J(l_{n})$. So, if we let $U_{j}=G_{j}(X)/\sigma_{j}^{\tau}$, then the previous two displays imply
$\mathbf{P}\big(\tilde{M}_{k_{n}}(X)\leq t\big)$ $\,\leq\,\mathbf{P}\Big(\max_{j\in J(l_{n})}\frac{G_{j}(X)}{\sigma_{j}}\leq c_{0}l_{n}^{(1-\tau)\beta}t\Big).$
Consequently, if we let $\omega$ be as defined in (24), and let $c_{2}=\frac{1}{c_{0}}\omega\sqrt{2(1-\omega)}$ in the definition (26) of $s_{2,n}$, then choosing $t=s_{2,n}$ in the previous display gives
$\mathbf{P}\Big(\tilde{M}_{k_{n}}(X)\leq s_{2,n}\Big)\ \leq\ \mathbf{P}\Big(\max_{j\in J(l_{n})}\frac{G_{j}(X)}{\sigma_{j}}\ \leq\omega\sqrt{2(1-\omega)\log(d_{n})}\Big).$
To bound the probability on the right, we apply Lemma 17 with $(l_{n},d_{n},\omega,\omega)$ playing the roles of $(d,k,a,b)$ in the statement of that result, which yields
$\mathbf{P}\Big(\tilde{M}_{k_{n}}(X)\leq s_{2,n}\Big)\ \lesssim\ d_{n}^{-(1-\omega)^{3}/\omega}(\log(d_{n}))^{\frac{1-\omega(2-\omega)-\omega}{2\omega}}.$
(Note that the conditions of Lemma 17 are applicable because $2\leq d_{n}\leq\frac{\omega^{2}}{4}{\tt{r}}(R(l_{n}))$ when $n$ is sufficiently large.) Furthermore, by Assumption 1((iv)), we have
$d_{n}\asymp{\tt{r}}(R(l_{n}))=\frac{l_{n}^{2}}{\|R(l_{n})\|_{F}^{2}}\ \gtrsim\ l_{n}^{\frac{1}{C}}\ \gtrsim\ n^{\omega}.$
Hence, there is a constant $c>0$ not depending on $n$ such that
$\begin{split}\mathbf{P}\Big(\tilde{M}_{k_{n}}(X)\leq s_{2,n}\Big)&\ \lesssim\ d_{n}^{-\frac{(1-\omega)^{3}}{\omega}}\log(d_{n})^{c}\\ &\ \lesssim\ n^{-(1-\omega)^{3}}\log(n)^{c}\\ &\ \lesssim\ n^{-1/2}\end{split}$
as needed. ∎
Lemma 2.
If the conditions of Theorem 1 hold, then there is a constant $c_{1}>0$ not depending on $n$ such that the following bound holds when $s_{1,n}=c_{1}k_{n}^{-\beta(1-\tau)/2}(2\vee\log(n))$,
$\mathbf{P}\big(B(s_{1,n})\big)\,\lesssim\,\textstyle\frac{1}{n}.$
Proof. Let $q=2\vee\log(n)$ and observe that
$\begin{split}\bigg\|\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\hat{Y}_{ij}}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}^{q}\ \leq\ \sum_{j\in J(k_{n})^{c}}\Bigg(2^{q}\bigg\|\frac{\sum_{i=1}^{n}\hat{Y}_{ij}-\mathbf{E}(\hat{Y}_{ij}|X^{\prime})}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}^{q}\ +\ 2^{q}\bigg|\frac{\sum_{i=1}^{n}\mathbf{E}(\hat{Y}_{ij}|X^{\prime})}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg|^{q}\Bigg)\end{split}$
holds almost surely. By Rosenthal’s inequality (Lemma 16), the following event holds almost surely,
$\begin{split}\bigg\|\frac{\sum_{i=1}^{n}\hat{Y}_{ij}-\mathbf{E}(\hat{Y}_{ij}|X^{\prime})}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}&\ \leq\ \frac{cq}{\hat{\sigma}_{j}^{\tau}}\max\bigg\{\sqrt{\textstyle\textup{var}(\hat{Y}_{1j}|X^{\prime})}\ ,n^{-1/2+1/q}\|\hat{Y}_{1j}\|_{L^{q}|X^{\prime}}\bigg\}\\[5.69046pt] &\ \leq\ \frac{cq}{\hat{\sigma}_{j}^{\tau}}\max\Big\{\sigma_{j}\ ,n^{1/q}\hat{\sigma}_{j}\Big\},\end{split}$
where the second step follows from $\textup{var}(\hat{Y}_{1j}|X^{\prime})\leq\mathbf{E}(X_{1j}^{2})\leq\sigma_{j}^{2}$ and $|\hat{Y}_{1j}|\leq n^{1/2}\hat{\sigma}_{j}$. Lemmas 13(iv) and 15 imply that there is an constant $c>0$ not depending on $n$ such that both of the bounds
$\max_{j\in J(k_{n})^{c}}\frac{\hat{\sigma}_{j}^{1-\tau}}{\sigma_{j}^{(1-\tau)/2}}\leq c\ \ \ \text{ and }\ \ \ \max_{j\in J(k_{n})^{c}}\frac{\sigma_{j}}{\hat{\sigma}_{j}^{\tau}\sigma_{j}^{(1-\tau)/2}}\leq c$
hold simultaneously with probability at least $1-cn^{-(2+\delta)}$ . Consequently, the bound
$\begin{split}\bigg\|\frac{\sum_{i=1}^{n}\hat{Y}_{ij}-\mathbf{E}(\hat{Y}_{ij}|X^{\prime})}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}&\leq cq\max\bigg\{\sigma_{j}^{(1-\tau)/2},\ \sigma_{j}^{(1-\tau)/2}n^{1/q}\bigg\}\\ &\leq cq\sigma_{j}^{(1-\tau)/2}\end{split}$
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(2+\delta)}$. Using Lemma 10 and similar reasoning, it can also be shown that
$\Big|\frac{\mathbf{E}(\hat{Y}_{ij}|X^{\prime})}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\Big|\leq c\sigma_{j}^{(1-\tau)/2}n^{-2}$
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(2+\delta)}$. Combining the last several steps and Assumption 1((iii)), we conclude that the bound
$\begin{split}\bigg\|\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\hat{Y}_{ij}}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}^{q}&\ \leq\ (cq)^{q}\sum_{j\geq k_{n}}j^{-q\beta(1-\tau)/2}\\[5.69046pt] &\ \leq\ c^{q}\frac{q^{q}}{(1-\tau)q\beta/2-1}k_{n}^{1-(1-\tau)q\beta/2}\end{split}$
holds with probability at least $1-cn^{-(2+\delta)}$, where in the second step we have used the fact that $q\beta(1-\tau)>2$ when $n$ is sufficiently large. Also, since $q\asymp\log(n)$, we have
$\Big(\frac{1}{(1-\tau)q\beta/2-1}\Big)^{1/q}\lesssim 1.$
Thus, there is a sufficiently large choice of $c_{1}>0$, such that if $s_{1,n}=c_{1}qk_{n}^{-\beta(1-\tau)/2}$, then the bound
$\mathbf{P}(B(s_{1,n})|X^{\prime})\,\leq\,\frac{1}{s_{1,n}^{q}}\bigg\|\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\hat{Y}_{ij}}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}^{q}\,\leq\,\textstyle e^{-q}\,\lesssim\,\textstyle\frac{1}{n}$
holds with probability at least $1-cn^{-(2+\delta)}$. This implies the stated result.∎
Proposition 4.
If the conditions of Theorem 1 hold, then
$d_{\textup{K}}\Big(\mathcal{L}(M_{k_{n}}(\hat{Y})),\mathcal{L}(M_{k_{n}}(Y))\Big)\lesssim n^{-\frac{1}{2}+\epsilon}.$
Proof. Using the decomposition (13) for the Kolmogorov metric, followed by the bound for anti-concentration probabilities in (14), we have
$d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(\hat{Y})),\mathcal{L}(M_{k_{n}}(Y)))\leq$ $\ \sup_{t\in\mathbb{R}}\mathbf{P}\big(|\tilde{M}_{k_{n}}(X)-t|\leq n^{-\frac{1}{2}+\frac{3\epsilon}{4}}\log(n)\big)$
+ $\ +\ 2d_{\textup{K}}(\mathcal{L}(\tilde{M}_{k_{n}}(X)),\mathcal{L}(M_{k_{n}}(Y)))$
$\ +\ \mathbf{P}\Big(\big|M_{k_{n}}(\hat{Y})-M_{k_{n}}(Y)\big|\geq n^{-\frac{1}{2}+\frac{3\epsilon}{4}}\log(n)\Big).$
For the first term on the right side, Nazarov’s inequality (Lemma 20) and Assumption 1((iii)) imply
$\begin{split}\sup_{t\in\mathbb{R}}\mathbf{P}\Big(|\tilde{M}_{k_{n}}(X)-t|\leq n^{-\frac{1}{2}+\frac{3\epsilon}{4}}\log(n)\Big)&\lesssim\,n^{-\frac{1}{2}+\frac{3\epsilon}{4}}\log(n)k_{n}^{\beta(1-\tau)}\sqrt{\log(k_{n})}\\ &\lesssim\ n^{-\frac{1}{2}+\epsilon}.\end{split}$ (28)
Next, it follows from Proposition 5 that
$d_{\textup{K}}\Big(\mathcal{L}(\tilde{M}_{k_{n}}(X)),\mathcal{L}(M_{k_{n}}(Y))\Big)\ \lesssim\ n^{-\frac{1}{2}+\epsilon}.$
Finally, Lemma 3 implies
$\mathbf{P}\Big(\big|M_{k_{n}}(\hat{Y})-M_{k_{n}}(Y)\big|\geq n^{-\frac{1}{2}+\frac{\epsilon}{2}}\log(n)\Big)\lesssim\ n^{-\frac{1}{2}+\epsilon},$
which completes the proof. ∎
Lemma 3.
If the conditions of Theorem 1 hold, then
$\mathbf{P}\Big(\big|M_{k_{n}}(\hat{Y})-M_{k_{n}}(Y)\big|\geq n^{-1/2+\epsilon/2}\log(n)\Big)\lesssim n^{-1/2+\epsilon}.$
Proof. First observe that $\big|M_{k_{n}}(\hat{Y})-M_{k_{n}}(Y)\big|$ can be bounded by
$\begin{split}&\max_{j\in J(k_{n})}\!\bigg|\frac{\sum_{i=1}^{n}\hat{Y}_{ij}-Y_{ij}}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg|+\max_{j\in J(k_{n})}\!\bigg|\frac{\sum_{i=1}^{n}Y_{ij}}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}-\frac{\sum_{i=1}^{n}Y_{ij}}{\sigma_{j}^{\tau}\sqrt{n}}\bigg|+\max_{j\in J(k_{n})}\!\frac{\sqrt{n}|\mathbf{E}(Y_{1j})|}{\sigma_{j}^{\tau}}.\end{split}$ (29)
The first term in the bound (29) is 0 with probability at least $1-\frac{ck_{n}}{n}$ by Lemma 4, and the (deterministic) third term is $\mathcal{O}(n^{-1})$ by Lemma 10.
It remains to handle the middle term in the bound (29), which satisfies
$\max_{j\in J(k_{n})}\big|\frac{\sum_{i=1}^{n}Y_{ij}}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}-\frac{\sum_{i=1}^{n}Y_{ij}}{\sigma_{j}^{\tau}\sqrt{n}}\big|\leq\max_{j\in J(k_{n})}\Big|\Big(\frac{{\sigma}_{j}}{\hat{\sigma}_{j}}\Big)^{\tau}-1\Big|\cdot\max_{j\in J(k_{n})}\bigg|\frac{\sum_{i=1}^{n}Y_{ij}}{\sqrt{n}\sigma_{j}^{\tau}}\bigg|.$ (30)
Since $|a^{\tau}-1|\leq|a^{2}-1|$ for any $a\geq 0$ and $\tau\in[0,1)$, the first factor on the right is of order $n^{-1/2+\epsilon/2}$ with probability at least $1-cn^{-(2+\delta)}$ by Lemma 13(iii).
Next we will show the second factor in the bound (30) is of order $\log(n)$ with probability at least $1-\frac{c}{n}$. Let $q=2\vee\log(n)$ and observe that under Assumption 1((iii)), Lemma 10 implies
$\bigg\|\max_{j\in J(k_{n})}\Big|\frac{\sum_{i=1}^{n}Y_{ij}}{\sqrt{n}\sigma_{j}^{\tau}}\Big|\bigg\|_{L^{q}}\ \lesssim\ \bigg\|\max_{j\in J(k_{n})}\Big|\sum_{i=1}^{n}\frac{Y_{ij}-\mathbf{E}(Y_{ij})}{\sqrt{n}\sigma_{j}^{\tau}}\Big|\bigg\|_{L^{q}}+\frac{1}{n}.$
Furthermore, Lemma 11 gives
$\bigg\|\max_{j\in J(k_{n})}\bigg|\sum_{i=1}^{n}\frac{Y_{ij}-\mathbf{E}(Y_{ij})}{\sqrt{n}\sigma_{j}^{\tau}}\bigg|\bigg\|_{L^{q}}^{q}$ $\leq\sum_{j\in J(k_{n})}\Big\|\sum_{i=1}^{n}\frac{Y_{ij}-\mathbf{E}(Y_{ij})}{\sqrt{n}\sigma_{j}^{\tau}}\Big\|_{L^{q}}^{q}$
$\leq c^{q}q^{q}\sum_{j\in J(k_{n})}\sigma_{j}^{(1-\tau)q}$
$\leq c^{q}q^{q}.$
Therefore, Chebyshev’s inequality implies that the second factor in the bound (30) is of order $\log(n)$ with probability $1-\frac{c}{n}$. ∎
Lemma 4.
If the conditions of Theorem 1 hold, then
$\mathbf{P}\Big(\max_{j\in J(k_{n})}\max_{1\leq i\leq n}|Y_{ij}-\hat{Y}_{ij}|>0\Big)\,\lesssim\,\frac{k_{n}}{n}.$
Proof. First observe that for each $i$ and $j$, the definitions of $Y_{ij}$ and $\hat{Y}_{ij}$ give
$\mathbf{P}\big(|Y_{ij}-\hat{Y}_{ij}|>0\big)\ \leq\ \mathbf{P}\Big(|X_{ij}|>\min\{t_{j},\hat{t}_{j}\}\Big).$
Recalling that $\hat{t}_{j}=\hat{\sigma}_{j}\sqrt{n}$ and $t_{j}=\sigma_{j}\sqrt{n}$, the events $\{\hat{t}_{j}>t_{j}/2\}$ for $j\in J(k_{n})$ occur simultaneously with probability at least $1-cn^{-(2+\delta)}$ by Lemma 13(iii). Combining this with a union bound over $i=1,\dots,n$, we have
$\mathbf{P}\Big(\max_{1\leq i\leq n}|Y_{ij}-\hat{Y}_{ij}|>0\Big)$ $\ \lesssim\ \sum_{i=1}^{n}\Big(\mathbf{P}\big(|X_{1j}|>\textstyle\frac{t_{j}}{2}\big)\ +\ n^{-(2+\delta)}\Big)$ (31)
$\ \lesssim\ \frac{n\,\mathbf{E}|X_{1j}|^{4}}{t_{j}^{4}}+\ n^{-(1+\delta)}$
$\ \lesssim\,\frac{1}{n},$
where the last step uses Assumption 1((i)). Finally, taking a union bound over $j\in J(k_{n})$,
$\mathbf{P}\Big(\max_{j\in J(k_{n})}\max_{1\leq i\leq n}|Y_{ij}-\hat{Y}_{ij}|>0\Big)\lesssim\frac{k_{n}}{n}$
as needed. ∎
Lemma 5.
If the conditions of Theorem 1 hold, then
$\mathbf{P}\Big(|M_{k_{n}}(Y)-M_{k_{n}}(X)|\geq n^{-\frac{1}{2}}\Big)\ \lesssim\ n^{-\frac{1}{2}+\epsilon},$
Proof. For each $i=1,\dots,n$ and $j\in J(k_{n})$, let
$\Delta_{ij}=\frac{1}{\sigma_{j}^{\tau}\sqrt{n}}\Big(Y_{ij}-\mathbf{E}(Y_{ij})-X_{ij}\Big),$
so that
$|M_{k_{n}}(Y)-M_{k_{n}}(X)|\ \leq\ \max_{j\in J(k_{n})}\Big|\sum_{i=1}^{n}\Delta_{ij}\Big|.$
Noting that $Y_{ij}$ and $X_{ij}$ only differ when $|X_{ij}|>t_{j}$, we have
$\mathbf{E}\big(|\Delta_{ij}|\big)$ $\ \lesssim\ \mathbf{E}\Big({\textstyle\frac{1}{\sqrt{n}\sigma_{j}^{\tau}}}|X_{ij}|1\{|X_{ij}|>t_{j}\}\Big)+\frac{1}{\sqrt{n}\sigma_{j}^{\tau}}\big|\mathbf{E}(Y_{ij})\big|$
$\ \lesssim\ \frac{1}{\sqrt{n}\sigma_{j}^{\tau}}\|X_{ij}\|_{L^{\frac{4}{1+\epsilon}}}\|1\{|X_{ij}|>t_{j}\}\|_{L^{\frac{4}{3-\epsilon}}}+n^{-2}$
$\ \lesssim\ n^{-2+\frac{\epsilon}{2}}.$
where we have used Hölder’s inequality and Lemma 10 in the first step, followed by Chebyshev’s inequality in bounding $\|1\{|X_{ij}|>t_{j}\}\|_{L^{\frac{4}{3-\epsilon}}}$. Therefore,
$\mathbf{P}\Big(\max_{j\in J(k_{n})}\big|\sum_{i=1}^{n}\Delta_{ij}\big|\geq n^{-\frac{1}{2}}\Big)\$$\ \leq\ \sum_{j\in J(k_{n})}\mathbf{P}\Big(\big|\sum_{i=1}^{n}\Delta_{ij}\big|\geq n^{-\frac{1}{2}}\Big)$
$\ \lesssim\ k_{n}n^{-\frac{1}{2}+\frac{\epsilon}{2}}$
$\ \lesssim\ n^{-\frac{1}{2}+\epsilon},$
as needed. ∎
Proposition 5.
If the conditions of Theorem 1 hold, then
$d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(Y)),\mathcal{L}(\tilde{M}_{k_{n}}(X)))\ \lesssim\ n^{-\frac{1}{2}+\epsilon}.$
Proof. First observe that
$d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(Y)),\mathcal{L}(\tilde{M}_{k_{n}}(X)))\ \leq\ d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(Y)),\mathcal{L}(M_{k_{n}}(X)))+d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(X)),\mathcal{L}(\tilde{M}_{k_{n}}(X))).$
The second term on the right side is of $n^{-1/2+\epsilon}$ by Lemma 6. Using the decomposition (13) for the Kolmogorov metric, the first term on the right side can be bounded by
$d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(Y)),\mathcal{L}(M_{k_{n}}(X)))\leq\sup_{t\in\mathbb{R}}\mathbf{P}(|M_{k_{n}}(X)-t|\leq n^{-\frac{1}{2}})+\mathbf{P}(|M_{k_{n}}(X)-M_{k_{n}}(Y)|>n^{-\frac{1}{2}}).$
Lemma 5 shows that
$\mathbf{P}(|M_{k_{n}}(X)-M_{k_{n}}(Y)|>n^{-\frac{1}{2}})\lesssim n^{-\frac{1}{2}+\epsilon}.$
Using the bound for anti-concentration probabilities in (14), we have
$\sup_{t\in\mathbb{R}}\mathbf{P}(|M_{k_{n}}(X)-t|\leq n^{-\frac{1}{2}})\leq$ $\ \sup_{t\in\mathbb{R}}\mathbf{P}(|\tilde{M}_{k_{n}}(X)-t|\leq n^{-\frac{1}{2}})$
$+2d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(X)),\mathcal{L}(\tilde{M}_{k_{n}}(X))).$
Nazarov’s inequality (Lemma 20) and Assumption 1((iii)) imply
$\begin{split}\sup_{t\in\mathbb{R}}\mathbf{P}(|\tilde{M}_{k_{n}}(X)-t|\leq n^{-\frac{1}{2}}))\lesssim\ n^{-\frac{1}{2}+\epsilon},\end{split}$ (32)
which completes the proof.
Lemma 6.
If the conditions of Theorem 1 hold, then
$d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(X)),\mathcal{L}(\tilde{M}_{k_{n}}(X)))\lesssim n^{-1/2+\epsilon}.$
Proof. For each $i=1,\dots,n$, let $X_{i}(k_{n})$ denote the vector in $\mathbb{R}^{k_{n}}$ corresponding to the coordinates of $X_{i}$ indexed by $J(k_{n})$, and let $\mathcal{R}_{t}=\prod_{j\in J(k_{n})}(-\infty,t\sigma_{j}^{\tau}]$, so that
$\mathbf{P}(M_{k_{n}}(X)\leq t)=\mathbf{P}\Big(\textstyle\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i}(k_{n})\in\mathcal{R}_{t}\Big).$
If the rank of the covariance matrix of $X_{1}(k_{n})$ is denoted by $r$, let $\Pi_{r}\in\mathbb{R}^{k_{n}\times r}$ be the matrix whose columns correspond to the leading $r$ eigenvectors of the covariance matrix of $X_{1}(k_{n})$. In particular, we have $X_{1}(k_{n})=\Pi_{r}\Pi_{r}^{\top}X_{1}(k_{n})$ almost surely, and it follows that
$\mathbf{P}(M_{k_{n}}(X)\leq t)=\mathbf{P}\Big(\textstyle\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Pi_{r}^{\top}X_{i}(k_{n})\in\Pi_{r}^{-1}(\mathcal{R}_{t})\Big),$
where $\Pi_{r}^{-1}(\mathcal{R}_{t})$ refers to the pre-image. Next, the definition of $\Pi_{r}$ ensures that the covariance matrix of $\Pi_{r}^{\top}X_{1}(k_{n})$, denoted by $\mathfrak{S}_{r}$, is invertible. So, if we define the random vector
$V_{i}=\mathfrak{S}_{r}^{-1/2}\Pi_{r}^{\top}X_{i}(k_{n})$
for each $i=1,\dots,n$ and the set $\mathcal{C}_{t}=\mathfrak{S}_{r}^{-1/2}\Pi_{r}^{-1}(\mathcal{R}_{t})$ for each $t\in\mathbb{R}$, then
$\mathbf{P}(M_{k_{n}}(X)\leq t)=\mathbf{P}\Big(\textstyle\frac{1}{\sqrt{n}}\sum_{i=1}^{n}V_{i}\in\mathcal{C}_{t}\Big).$
It can also be shown by similar reasoning that
$\mathbf{P}(\tilde{M}_{k_{n}}(X)\leq t)=\gamma_{r}(\mathcal{C}_{t}),$
where $\gamma_{r}$ denotes the standard Gaussian distribution on $\mathbb{R}^{r}$. Due to the fact that the i.i.d. random vectors $V_{1},\dots,V_{n}$ are centered and isotropic, Bentkus’ multivariate Berry-Esseen theorem (Lemma 19) ensures there is an absolute constant $c>0$ such that
$\begin{split}d_{\textup{K}}(\mathcal{L}(M_{k_{n}}(X)),\mathcal{L}(\tilde{M}_{k_{n}}(X)))&\leq\ \sup_{\mathcal{C}\in\mathscr{C}}\Big|\mathbf{P}\Big(\textstyle\frac{1}{\sqrt{n}}\sum_{i=1}^{n}V_{i}\in\mathcal{C}\Big)-\gamma_{r}(\mathcal{C})\Big|\\[5.69046pt] &\leq\ \displaystyle\frac{cr^{1/4}\mathbf{E}\|V_{1}\|_{2}^{3}}{\sqrt{n}},\end{split}$ (33)
where $\mathscr{C}$ denotes the collection of all Borel convex subsets of $\mathbb{R}^{r}$. To bound the third moment on the right side, observe that
$\mathbf{E}\|V_{1}\|_{2}^{3}$ $\ =\ \Big\|\textstyle\sum_{j=1}^{r}V_{1j}^{2}\Big\|_{L^{3/2}}^{3/2}\ \leq\ \Big(\textstyle\sum_{j=1}^{r}\|V_{1j}\|_{L^{3}}^{2}\Big)^{3/2}.$
Since $V_{1j}$ can be expressed as $\langle v,X_{1}\rangle$ for some vector $v\in\mathbb{R}^{p}$, Assumption 1((i)) implies $\|V_{1j}\|_{L^{3}}^{2}\lesssim\textup{var}(V_{1j})=1$ for each $j=1,\dots,r$. Thus $\mathbf{E}\|V_{1}\|_{2}^{3}\lesssim r^{3/2}\leq k_{n}^{3/2}$, and combining with (33) completes the proof.∎
Appendix D Gaussian comparison
Recall that $M_{k_{n}}^{*}(X)=\max_{j\in J(k_{n})}\frac{1}{\sqrt{n}\sigma_{j}^{\tau}}\sum_{i=1}^{n}\xi_{i}(X_{ij}-\bar{X}_{j})$ from (12).
Proposition 6.
Suppose the conditions of Theorem 1 hold. Then, there is a constant $c>0$ not depending on $n$ such that the event
$d_{\textup{K}}\Big(\mathcal{L}(\tilde{M}_{k_{n}}(X))\,,\,\mathcal{L}(M_{k_{n}}^{*}(X)|X)\Big)\leq cn^{-\frac{1}{2}+\epsilon}$
holds with probability at least $1-cn^{-\delta/4}$.
Proof. Let $V_{1},\dots,V_{n}\in\mathbb{R}^{r}$ be as in the proof of Lemma 6 and put $\bar{V}=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. The reasoning used in the proof of Lemma 6 shows that for any $t\in\mathbb{R}$, there is a convex Borel set $\mathcal{C}_{t}\subset\mathbb{R}^{r}$ such that
$\mathbf{P}\big(M_{k_{n}}^{*}(X)\leq t|X\big)\ =\ \mathbf{P}\Big(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\xi_{i}(V_{i}-\bar{V})\in\mathcal{C}_{t}\Big|X\Big),$
and also, that
$\mathbf{P}(\tilde{M}_{k_{n}}(X)\leq t)=\gamma_{r}(\mathcal{C}_{t}),$
where $\gamma_{r}$ is the standard Gaussian distribution on $\mathbb{R}^{r}$. Next, define the sample covariance matrix
$W_{r}=\frac{1}{n}\sum_{i=1}^{n}(V_{i}-\bar{V})(V_{i}-\bar{V})^{\top}$
and observe that Lemma 18 gives the following almost-sure bound,
$\sup_{t\in\mathbb{R}}\Big|\mathbf{P}\big(M_{k_{n}}^{*}(X)\leq t\big|X\big)\ -\ \mathbf{P}\big(\tilde{M}_{k_{n}}(X)\leq t\big)\Big|\ \leq\ 2\big\|W_{r}-I_{r}\big\|_{F},$
where $I_{r}$ denotes the identity matrix of size $r\times r$. To handle the Frobenius norm, Assumption 1((i)) implies
$\max_{1\leq j,j^{\prime}\leq r}\mathbf{E}\Big|e_{j}^{\top}(V_{1}V_{1}^{\top}-I_{r})e_{j^{\prime}}\Big|^{\frac{4+\delta}{2}}\lesssim 1,$
$\max_{1\leq j\leq r}\mathbf{E}\big|e_{j}^{\top}V_{1}\big|^{4+\delta}\lesssim 1.$
Consequently, the Fuk-Nagaev inequality (Lemma 21) with the choices $q=(4+\delta)/2$ and $q=4+\delta$ in the notation of that result ensures that for each $j,j^{\prime}=1,\ldots,r$,
$\mathbf{P}\bigg(\Big|\frac{1}{n}\sum_{i=1}^{n}e_{j}^{\top}(V_{i}V_{i}^{\top}-I_{r})e_{j^{\prime}}\Big|\geq\textstyle n^{-1/2+\epsilon/2}\bigg)\ \lesssim\ n^{-(\delta/4+\epsilon)},$
$\mathbf{P}\Big(\big|e_{j}^{\top}\bar{V}\big|\geq n^{-1/4+\epsilon/4}\Big)\leq n^{-(2+3\delta/4+\epsilon)}.$
So, using the identity
$W_{r}-I_{r}=\Big(\frac{1}{n}\sum_{i=1}^{n}V_{i}V_{i}^{\top}-I_{r}\Big)-\bar{V}\bar{V}^{\top}$
it is straightforward to check that the event
$\big\|W_{r}-I_{r}\big\|_{F}\geq 2k_{n}n^{-1/2+\epsilon/2}$
holds with probability at most of order $k_{n}^{2}n^{-(\delta/4+\epsilon)}\lesssim n^{-\delta/4}$, which leads to the stated result.∎
Appendix E Bootstrap approximation
Proposition 7.
Suppose the conditions of Theorem 1 hold. Then, there is a constant $c>0$ not depending on $n$ such that the event
$d_{\textup{K}}\Big(\mathcal{L}(M_{k_{n}}^{*}(X)|X),\mathcal{L}(M_{p}^{*}(\hat{Z})|X))\Big)\leq cn^{-\frac{1}{2}+\epsilon}$
holds with probability at least $1-cn^{-\delta/4}$.
Proof. Consider the inequality
$d_{\textup{K}}\Big(\mathcal{L}(M_{k_{n}}^{*}(X)|X),\mathcal{L}(M_{p}^{*}(\hat{Z})|X))\Big)\leq\mathrm{I}^{\prime}_{n}+\mathrm{II}^{\prime}_{n},$
where we define
$\mathrm{I}^{\prime}_{n}$ $=d_{\textup{K}}\Big(\mathcal{L}(M_{k_{n}}^{*}(X)|X),\mathcal{L}(M_{k_{n}}^{*}(\hat{Z})|X)\Big),$
$\mathrm{II}^{\prime}_{n}$ $=d_{\textup{K}}\Big(\mathcal{L}(M_{k_{n}}^{*}(\hat{Z})|X),\mathcal{L}(M_{p}^{*}(\hat{Z})|X)\Big).$
Both $\mathrm{I}^{\prime}_{n}$ and $\mathrm{II}^{\prime}_{n}$ are of order at most $n^{-1/2+\epsilon}$ with probability at least $1-cn^{-\delta/4}$, as shown in Lemma 7 and Proposition 8.∎
Lemma 7.
Suppose the conditions of Theorem 1 hold. Then, there is a constant $c>0$ not depending on $n$ such that the event
$d_{\textup{K}}\Big(\mathcal{L}(M_{k_{n}}^{*}(X)|X),\mathcal{L}(M_{k_{n}}^{*}(\hat{Z})|X)\Big)\leq cn^{-1/2+\epsilon}$
holds with probability at least $1-cn^{-\delta/4}$.
Proof. The coupling and anti-concentration decomposition in (13) shows that for any $\eta>0$, we have
$d_{\textup{K}}\Big(\mathcal{L}\big(M_{k_{n}}^{*}(X)|X\big),\mathcal{L}\big(M_{k_{n}}^{*}(\hat{Z})|X\big)\Big)\ \leq$ $\sup_{t\in\mathbb{R}}\mathbf{P}\Big(|M_{k_{n}}^{*}(X)-t|\leq\eta\Big|X\Big)$ (34)
$+\mathbf{P}\Big(\big|M_{k_{n}}^{*}(X)-M_{k_{n}}^{*}(\hat{Z})\big|\geq\eta\Big|X\Big).$
We will take $\eta=c\log(n)n^{-\frac{1}{2}+\frac{3\epsilon}{4}}$ for some constant $c>0$ not depending on $n$. Using the generic bound for anti-concentration probabilities in (14), the first term on the right side of (34) is upper bounded by
$\sup_{t\in\mathbb{R}}\mathbf{P}\big(|\tilde{M}_{k_{n}}(X)-t|\leq\eta\big)+\ 2d_{\textup{K}}\Big(\mathcal{L}\big(\tilde{M}_{k_{n}}(X)\big),\mathcal{L}\big(M_{k_{n}}^{*}(X)|X\big)\Big),$
which is at most of order $n^{-1/2+\epsilon}$ with probability at least $1-cn^{-\delta/4}$, due to (28) and Proposition 6.
To address the coupling term in (34), observe that for any $q\geq 2$, the basic inequalities (15), (17) and (18) ensure there is a constant $c>0$ not depending on $n$ such that
$\big\|M_{k_{n}}^{*}(X)-M_{k_{n}}^{*}(\hat{Z})\big\|_{L^{q}|X}\ \leq\ c\sqrt{q}k_{n}^{1/q}\max_{j\in J(k_{n})}\bigg(\frac{1}{n}\sum_{i=1}^{n}\Big(\textstyle\frac{X_{ij}-\bar{X}_{j}}{\sigma_{j}^{\tau}}-\frac{\hat{Z}_{ij}-\hat{Z}_{j}}{\hat{\sigma}_{j}^{\tau}}\Big)^{2}\bigg)^{1/2}.$
To decompose this bound, define the random variables
$T_{1}$ $=\displaystyle\max_{j\in J(k_{n})}\bigg(\frac{1}{n}\sum_{i=1}^{n}\Big(\textstyle\frac{X_{ij}-\hat{Z}_{ij}}{\sigma_{j}^{\tau}}\Big)^{2}\bigg)^{1/2},$
$T_{2}$ $=\displaystyle\max_{j\in J(k_{n})}\Big|\frac{1}{\sigma_{j}^{\tau}}-\frac{1}{\hat{\sigma}_{j}^{\tau}}\Big|\bigg(\frac{1}{n}\sum_{i=1}^{n}\hat{Z}_{ij}^{2}\bigg)^{1/2}.$
Using Jensen’s inequality for sample averages, it follows that
$\big\|M_{k_{n}}^{*}(X)-M_{k_{n}}^{*}(\hat{Z})\big\|_{L^{q}|X}\ \leq\ c\sqrt{q}k_{n}^{1/q}(T_{1}+T_{2}).$
With regard to $T_{1}$, note that
$\begin{split}|X_{ij}-\hat{Z}_{ij}|&\ \leq\ |\tilde{X}_{j}|1\{|X_{ij}-\tilde{X}_{j}|\leq\hat{t}_{j}\}\ +\ (|X_{ij}|+\hat{t}_{j})1\big\{|X_{ij}-\tilde{X}_{j}|>\hat{t}_{j}\big\}.\end{split}$
Also, by Lemma 13(iii), the events
$\max_{j\in J(k_{n})}\frac{\hat{t}_{j}}{t_{j}}\ \leq\ 2\ \ \ \ \ \text{ and }\ \ \ \ \ \min_{j\in J(k_{n})}\frac{\hat{t}_{j}}{t_{j}}\ \geq\ \textstyle\frac{1}{2}$
hold simultaneously with probability at least $1-cn^{-(2+\delta)}$ for some constant $c>0$ not depending on $n$. Based on this and $\min_{j\in J(k_{n})}\sigma_{j}^{\tau}\gtrsim k_{n}^{-\beta\tau}$ under Assumption 1((iii)), the bound
$\Big(\textstyle\frac{X_{ij}-\hat{Z}_{ij}}{\sigma_{j}^{\tau}}\Big)^{2}\ \leq\ ck_{n}^{2\beta\tau}\tilde{X}_{j}^{2}\ +\ ck_{n}^{2\beta\tau}(X_{ij}^{2}+t_{j}^{2})\Big(1\big\{|X_{ij}|>\textstyle\frac{t_{j}}{4}\big\}+1\big\{|\tilde{X}_{j}|>\textstyle\frac{t_{j}}{4}\big\}\Big)$
holds simultaneously for all $i\in\{1,\dots,n\}$ and $j\in J(k_{n})$ with probability at least $1-cn^{-(2+\delta)}$. Therefore, the bound
$T_{1}\leq ck_{n}^{\beta\tau}\max_{j\in J(k_{n})}\Bigg(|\tilde{X}_{j}|+\bigg(\frac{1}{n}\sum_{i=1}^{n}\Big(X_{ij}^{2}+t_{j}^{2}\Big)1\{|X_{ij}|>\textstyle\frac{t_{j}}{4}\}\bigg)^{\frac{1}{2}}+1\{|\tilde{X}_{j}|>\textstyle\frac{t_{j}}{4}\}\bigg(\displaystyle\frac{1}{n}\sum_{i=1}^{n}X_{ij}^{2}+t_{j}^{2}\bigg)^{\frac{1}{2}}\Bigg)$
holds with the same probability. The terms on the right side are handled as follows. First, $\max_{j\in J(k_{n})}|\tilde{X}_{j}|$ is of order $n^{-1/2+\epsilon/2}$ with probability at least $1-cn^{-(2+\delta)}$ by Assumption 1((iii)) and Lemma 12(iii). The indicators $1\{|X_{ij}|\geq\frac{t_{j}}{4}\}$ and $1\{|\tilde{X}_{j}|\geq\frac{t_{j}}{4}\}$ are 0 with probability at least $1-cn^{-2}$ due to the argument associated with the bounds in (31), as well as Lemma 12(iii) and the fact that $\sigma_{j}n^{-\frac{1}{2}+\frac{\epsilon}{2}}\lesssim\frac{t_{j}}{4}$. After taking a union bound over $i\in\{1,\dots,n\}$ and $j\in J(k_{n})$, the event
$T_{1}$ $\leq ck_{n}^{\beta\tau}n^{-1/2+\epsilon/2}\leq cn^{-1/2+3\epsilon/4}$
holds with probability at least $1-ck_{n}/n$.
Turning our attention to $T_{2}$, Lemma 13(iii) implies that the bound
$\max_{j\in J(k_{n})}\Big|\frac{1}{\sigma_{j}^{\tau}}-\frac{1}{\hat{\sigma}_{j}^{\tau}}\Big|\ \leq ck_{n}^{\beta\tau}n^{-1/2+\epsilon/2}\leq cn^{-1/2+3\epsilon/4},$
holds with probability at least $1-cn^{-(2+\delta)}$. Also, it will be shown in Equation (36) that the bound
$\max_{j\in J(k_{n})}\Big(\frac{1}{n}\sum_{i=1}^{n}\hat{Z}_{ij}^{2}\Big)^{1/2}\leq c\log(n)$
holds with probability at least $1-ck_{n}/n$. Combining the last two steps shows that $T_{2}$ is of order $\log(n)n^{-1/2+3\epsilon/4}$ with the same probability. ∎
Proposition 8.
Suppose the conditions of Theorem 1 hold. Then, there is a constant $c>0$ not depending on $n$ such that the event
$d_{\textup{K}}(\mathcal{L}(M_{k_{n}}^{*}(\hat{Z})|X),\mathcal{L}(M_{p}^{*}(\hat{Z})|X))\leq cn^{-\frac{1}{2}+\epsilon},$
holds with probability at least $1-cn^{-\delta/4}$.
[…]
$A^{\prime}(t)=\Big\{\max_{j\in J(k_{n})}\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}\leq t\Big\}\ \ \ \text{and}\ \ \ B^{\prime}(t)=\Big\{\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}>t\Big\}.$
Using the argument in the proof of Proposition 3, it can be shown that the following bound holds almost surely for any real numbers $s_{1,n}^{\prime}\leq s_{2,n}^{\prime}$,
$d_{\textup{K}}\Big(\mathcal{L}(M_{k_{n}}^{*}(\hat{Z})|X),\mathcal{L}(M_{p}^{*}(\hat{Z})|X)\Big)\ \leq\ \mathbf{P}\big(A^{\prime}(s_{2,n}^{\prime})|X\big)\,+\,\mathbf{P}\big(B^{\prime}(s_{1,n}^{\prime})|X\big).$
If we choose
$s_{1,n}^{\prime}=c_{1}^{\prime}k_{n}^{-\beta(1-\tau)/4}\log(n)^{3/2},\quad s_{2,n}^{\prime}=c_{2}^{\prime}l_{n}^{-\beta(1-\tau)}\sqrt{\log(d_{n})},$
where $c_{1}^{\prime},c_{2}^{\prime}>0$ are constants not depending on $n$, then $s_{1,n}^{\prime}\leq s_{2,n}^{\prime}$ holds when $n$ is sufficiently large (regardless of the particular values of $c_{1}^{\prime}$ and $c_{2}^{\prime}$). Recall also that $d_{n}$ is defined in (25). Lemma 8 shows that there is a choice of $c_{1}^{\prime}$ such that random variable $\mathbf{P}\big(B^{\prime}(s_{1,n}^{\prime})|X\big)$ is at most $n^{-1}$ with probability at least $1-cn^{-(1+\delta)}$. To deal with and $\mathbf{P}\big(A^{\prime}(s_{2,n}^{\prime})|X\big)$, notice that
$\mathbf{P}\big(A^{\prime}(s_{2,n}^{\prime})|X\big)\leq\mathbf{P}(\tilde{M}_{k_{n}}(X)\leq s_{2,n}^{\prime})\ +\ d_{\textup{K}}(\mathcal{L}(\tilde{M}_{k_{n}}(X),\mathcal{L}(M_{k_{n}}^{*}(\hat{Z})|X)).$
Lemma 1 shows there is a choice of $c_{2}^{\prime}$ such that the first term on the right hand side is of order $n^{-1/2}$. Finally, the second term is of order $n^{-1/2+\epsilon}$ with probability at least $1-cn^{-\delta/4}$ by Proposition 6 and Lemma 7. ∎
Lemma 8.
If the conditions of Theorem 1 hold, then there are constants $c,c_{1}^{\prime}>0$ not depending on $n$ such that the following event holds with probability at least $1-cn^{-(1+\delta)}$ when $s_{1,n}^{\prime}=c_{1}^{\prime}k_{n}^{-\beta(1-\tau)/4}\log(n)^{3/2}$,
$\mathbf{P}\big(B^{\prime}(s_{1,n}^{\prime})|X\big)\leq\frac{1}{n}.$
Proof.
Notice that
$\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}\leq\max_{j\in J(k_{n})^{c}}\frac{\sigma_{j}^{(\tau+1)/2}}{\hat{\sigma}_{j}^{\tau}}\ \cdot\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}\sigma_{j}^{(\tau+1)/2}}.$
The first factor on the right side is of order 1 with probability at least $1-cn^{-(2+\delta)}$ by Lemma 15. To handle the second factor, let $q=\max\{2,(1+\delta)\log(n)\}$. The idea of the rest of the proof is to construct a number $b_{n}$ such that the following event holds with high probability for every realization of the data,
$\Big\|\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}\sigma_{j}^{(\tau+1)/2}}\Big\|_{L^{q}|X}\leq b_{n}.$
This will lead to the statement of the lemma via Chebyshev’s inequality, because it will turn out that $s_{1,n}^{\prime}\asymp b_{n}$. To construct $b_{n}$, first observe that
$\Big\|\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}\sigma_{j}^{(\tau+1)/2}}\Big\|_{L^{q}|X}^{q}$ $\leq\sum_{j\in J(k_{n})^{c}}\sigma_{j}^{-q(\tau+1)/2}\Big\|\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}}\Big\|_{L^{q}|X}^{q}$
$\leq c^{q}q^{3q/2}\sum_{j\in J(k_{n})^{c}}\sigma_{j}^{(1-\tau)q/4}$
$\leq c^{q}\frac{q^{3q/2}}{(1-\tau)q\beta/4-1}k_{n}^{1-(1-\tau)q\beta/4}$
where the second inequality holds with probability at least $1-cn^{-(1+\delta)}$ by Lemma 9, and the third inequality follows from the fact that $q\beta(1-\tau)/4>1$ when $n$ is sufficiently large. Since $q\asymp\log(n)$, we have
$\Big(\frac{1}{(1-\tau)q\beta/4-1}k_{n}\Big)^{1/q}\lesssim 1,$
and so the event
$\bigg\|\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}\sigma_{j}^{(\tau+1)/2}}\bigg\|_{L^{q}|X}\leq cq^{3/2}k_{n}^{-(1-\tau)\beta/4}$
holds with probability at least $1-cn^{-(1+\delta)}$. Thus, we may take $b_{n}$ to be of the form $b_{n}=cq^{3/2}k_{n}^{-(1-\tau)\beta/4}$, and there is a choice of $c_{1}^{\prime}$ such that the stated result holds. ∎
Lemma 9.
Let $q=\max\{2,(1+\delta)\log(n)\}$ and suppose the conditions of Theorem 1 hold. Then, there is a constant $c>0$ not depending on $n$ such that the bound
$\bigg\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})\bigg\|_{L^{q}|X}\leq cq^{3/2}\sigma_{j}^{\frac{\tau+3}{4}}$
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(1+\delta)}$.
Proof. The $L^{q}$-norm bound for centered Gaussian random variables in (15) gives
$\bigg\|\frac{\sum_{i=1}^{n}\xi_{i}(\hat{Z}_{ij}-\hat{Z}_{j})}{\sqrt{n}}\bigg\|_{L^{q}|X}\ \leq\ c\sqrt{q}\bigg(\frac{1}{n}\sum_{i=1}^{n}\hat{Z}_{ij}^{2}\bigg)^{1/2}.$ (35)
To develop a high-probability bound for the right side, note that we have $|\hat{Z}_{ij}|\leq|\hat{Y}_{ij}|+|\tilde{X}_{j}|$ for any fixed $j$, and so
$\frac{1}{n}\sum_{i=1}^{n}\hat{Z}_{ij}^{2}\ \leq\ \frac{2}{n}\sum_{i=1}^{n}\hat{Y}_{ij}^{2}\ +\ 2|\tilde{X}_{j}|^{2}.$
Here, we apply Lemma 12(iv) with $\theta=(1-\tau)/4$ in the notation used there. This implies there is a constant $c>0$ not depending on $n$ such that the bound
$|\tilde{X}_{j}|^{2}\leq c\sigma_{j}^{\frac{\tau+3}{2}}$
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(1+\delta)}$. Next, Rosenthal’s inequality for non-negative random variables (Lemma 16) gives
$\begin{split}\bigg\|\frac{1}{n}\sum_{i=1}^{n}\hat{Y}_{ij}^{2}\bigg\|_{L^{q}|X^{\prime}}&\ \leq\ cq\max\Big\{\|\hat{Y}_{1j}^{2}\|_{L^{1}|X^{\prime}}\,,\,n^{-1+1/q}\|\hat{Y}_{1j}^{2}\|_{L^{q}|X^{\prime}}\Big\}\\[2.84544pt] &\ \leq\ cq\max\Big\{\|X_{1j}^{2}\|_{L^{1}}\,,\,n^{-1+1/q}\hat{t}_{j}^{2}\Big\}\\[2.84544pt] &\ \leq\ cq(\sigma_{j}^{2}+\hat{\sigma}_{j}^{2}),\end{split}$
where the second step uses $|\hat{Y}_{1j}|\leq|X_{1j}|\wedge\hat{t}_{j}$. Applying Chebyshev’s inequality conditionally on $X^{\prime}$ shows that the event
$\mathbf{P}\Big(\frac{1}{n}\sum_{i=1}^{n}\hat{Y}_{ij}^{2}\ \geq\ ceq(\sigma_{j}^{2}+\hat{\sigma}_{j}^{2})\sigma_{j}^{\frac{\tau-1}{4}}\Big|X^{\prime}\Big)\ \leq\ e^{-q}\sigma_{j}^{\frac{(1-\tau)q}{4}}\ \leq cn^{-(1+\delta)}\sigma_{j}^{\frac{(1-\tau)q}{4}}$
holds with probability 1, and thus the unconditional version of the left hand side is also at most $cn^{-(1+\delta)}\sigma_{j}^{\frac{(1-\tau)q}{4}}$. Consequently, the event
$\frac{1}{n}\sum_{i=1}^{n}\hat{Y}_{ij}^{2}\ \leq\ ceq(\sigma_{j}^{2}+\hat{\sigma}_{j}^{2})\sigma_{j}^{\frac{\tau-1}{4}}$
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(1+\delta)}$ since
$\sum_{j\in J(k_{n})^{c}}j^{-\frac{(1-\tau)q\beta}{4}}\lesssim k_{n}^{-\frac{(1-\tau)q\beta}{4}+1}\lesssim 1.$
Finally, Lemma 13(iv) implies there is a constant $c>0$ not depending on $n$ that the bound
$\hat{\sigma}_{j}^{2}\leq c\sigma_{j}^{\frac{\tau+7}{4}}$
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(2+\delta)}$. Combining results above, we have that the bound
$\frac{1}{n}\sum_{i=1}^{n}\hat{Z}_{ij}^{2}\ \leq\ cq\sigma_{j}^{\frac{\tau+3}{2}}$ (36)
holds simultaneously over $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(1+\delta)}$, which completes the proof.∎
Lemma 10.
If the conditions of Theorem 1 hold, then there is a constant $c>0$ not depending on $n$ such that the following bounds hold for all $j,k\in\{1,\dots,p\}$,
$|\mathbf{E}(Y_{1j})|$ $\,\leq\,c\sigma_{j}n^{-\frac{3}{2}}$
$|\mathbf{E}(\hat{Y}_{1j}|X^{\prime})|$ $\,\leq\,c\sigma_{j}^{4}\hat{\sigma}_{j}^{-3}n^{-\frac{3}{2}}\ \ \textup{ (almost surely)}$
$|\textup{cov}(Y_{1j},Y_{1k})-\textup{cov}(X_{1j},X_{1k})|$ $\,\leq\,c\sigma_{j}\sigma_{k}n^{-1}.$
Proof. We may assume without loss of generality that $\mathbf{E}(X_{1j})=0$ for all $j\in\{1,\dots,p\}$. Observe that Assumption 1((i)) gives
$|\mathbf{E}(Y_{1j})|$ $\leq\mathbf{E}(|X_{1j}|1\{|X_{1j}|\geq t_{j}\})$
$\leq\|X_{1j}\|_{L^{4}}\|1\{|X_{1j}|\geq t_{j}\}\|_{L^{4/3}}$
$\leq c\sigma_{j}\Big(\textstyle\frac{\|X_{1j}\|_{L^{4}}^{4}}{t_{j}^{4}}\Big)^{3/4}$
$\leq c\sigma_{j}n^{-\frac{3}{2}}.$
Second, the stated bound on $|\mathbf{E}(\hat{Y}_{1j}|X^{\prime})|$ can also be obtained from essentially the same argument. Third, to bound the difference between the covariances, we use the fact that $|Y_{1j}Y_{1k}-X_{1j}X_{1k}|$ vanishes on the intersection of the events $\{|X_{1j}|\leq t_{j}\}$ and $\{|X_{1k}|\leq t_{k}\}$, and otherwise it is at most $2|X_{1j}X_{1k}|$. Therefore, we have
$\big|\mathbf{E}(Y_{1j}Y_{1k})-\mathbf{E}(X_{1j}X_{1k})\big|$ $\ \leq\ 2\mathbf{E}\big(|X_{1j}X_{1k}|1\{|X_{1j}|\geq t_{j}\}\big)+2\mathbf{E}\big(|X_{1j}X_{1k}|1\{|X_{1k}|\geq t_{k}\}\big).$
The two terms on the right hand side can be handled via
$\mathbf{E}\big(|X_{1j}X_{1k}|1\{|X_{1j}|\geq t_{j}\}\big)$ $\leq\|X_{1j}\|_{L^{4}}\|X_{1k}\|_{L^{4}}\|1\{|X_{1j}|\geq t_{j}\}\big\|_{L^{2}}$
$\leq c\sigma_{j}\sigma_{k}\Big(\textstyle\frac{\|X_{1j}\|_{L^{4}}^{4}}{t_{j}^{4}}\Big)^{1/2}$
$\leq\frac{c\sigma_{j}\sigma_{k}}{n},$
which yields the stated result. ∎
Lemma 11.
If the conditions of Theorem 1 hold and $q=\max\{\log(n),2\}$, then
$\bigg\|\sum_{i=1}^{n}\frac{Y_{ij}-\mathbf{E}(Y_{ij})}{\sqrt{n}\sigma_{j}}\bigg\|_{L^{q}}\lesssim q.$
Proof.
Due to Rosenthal’s inequality (Lemma 16), we have
$\bigg\|\sum_{i=1}^{n}\frac{Y_{ij}-\mathbf{E}(Y_{ij})}{\sqrt{n}\sigma_{j}}\bigg\|_{L^{q}}$ $\lesssim q\max\Bigg\{\bigg\|\sum_{i=1}^{n}\frac{Y_{ij}-\mathbf{E}(Y_{ij})}{\sqrt{n}\sigma_{j}}\bigg\|_{L^{2}},\bigg(\sum_{i=1}^{n}\bigg\|\frac{Y_{ij}-\mathbf{E}(Y_{ij})}{\sqrt{n}\sigma_{j}}\bigg\|_{L^{q}}^{q}\bigg)^{1/q}\Bigg\}$
$\lesssim q\max\big\{1,n^{1/q}\big\}$
$\lesssim q,$
where the second step uses the almost-sure bound $|Y_{ij}|\leq\sqrt{n}\sigma_{j}$, as well as Lemma 10 to relate the variance of $Y_{ij}$ with $\sigma_{j}$. ∎
Appendix F Results on median-of-means estimators
The results in this section will continue to follow the convention that $\max_{1\leq j\leq p}\sigma_{j}^{2}=1$, as discussed on p.B. However, to make the results easier to interpret, we will state them so that they explicitly account for the coordinate-wise means $\mu_{j}=\mathbf{E}(X_{1j})$, $j=1,\dots,p$ (even though these parameters may be assumed to be zero without loss of generality in proving Theorem 1).
Lemma 12.
Fix any constant $\theta\in(0,1)$ and suppose that the conditions of Theorem 1 hold. Then, there is a constant $c\geq 1$ not depending on $n$, such that for any $j\in\{1,\dots,p\}$, the median-of-means estimator $\tilde{X}_{j}$ with $b_{n}\asymp\log(n)$ blocks satisfies
$\mathbf{P}\Big(|\tilde{X}_{j}-\mu_{j}|\geq\sigma_{j}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ \Big(\frac{c}{n}\Big)^{\frac{\epsilon}{c}b_{n}}$ (i)
$\mathbf{P}\Big(|\tilde{X}_{j}-\mu_{j}|\geq\sigma_{j}^{1-\theta}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ (c\sigma_{j})^{\frac{\theta}{c}b_{n}}.$ (ii)
Furthermore, we have
$\sum_{j\in J(k_{n})}\mathbf{P}\Big(|\tilde{X}_{j}-\mu_{j}|\geq\sigma_{j}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ n^{-(2+\delta)}$ (iii)
$\sum_{j=1}^{p}\mathbf{P}\Big(|\tilde{X}_{j}-\mu_{j}|\geq C\sigma_{j}^{1-\theta}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ k_{n}^{-\log(n)/c}.$ (iv)
Proof. Recall the notation $\bar{X}_{j}(l)=\frac{1}{\ell_{n}}\sum_{i\in\mathcal{B}_{l}}X_{ij}$ where $l=1,\dots,b_{n}$. Fix $t>0$ and let $\xi_{jl}=1\big\{|\bar{X}_{j}(l)-\mu_{j}|\geq t\big\}$. Since the event $\{|\tilde{X}_{j}(l)-\mu_{j}|\geq t\}$ can only occur if at least half of the random variables $\xi_{j1},\dots,\xi_{jb_{n}}$ are 1, we must have
$\mathbf{P}\big(|\tilde{X}_{j}-\mu_{j}|\geq t\big)$ $\ \leq\ \mathbf{P}\bigg(\frac{1}{b_{n}}\sum_{l=1}^{b_{n}}\xi_{jl}\ \geq\frac{1}{2}\bigg).$
Next, applying Kiefer’s inequality (Lemma 22) to the right side gives
$\mathbf{P}(|\tilde{X}_{j}-\mu_{j}|\geq t)\ \lesssim\ \big(e\mathbf{E}(\xi_{1j})\big)^{b_{n}(\frac{1}{2}-\mathbf{E}(\xi_{j1}))^{2}}.$ (37)
Furthermore, by Chebyshev’s inequality, $\mathbf{E}(\xi_{jl})\lesssim\ \frac{\sigma_{j}^{2}}{\ell_{n}t^{2}}$, and so if we take $t=\sigma_{j}n^{-1/2+\epsilon/2}$, then
$\begin{split}\mathbf{E}(\xi_{jl})\ \lesssim\ \frac{n^{1-\epsilon}}{\ell_{n}}\ \lesssim\ n^{-\epsilon/2},\end{split}$
where the last step uses $n/\ell_{n}\asymp m_{n}/\ell_{n}=b_{n}\asymp\log(n)$. Thus, combining this bound on $\mathbf{E}(\xi_{jl})$ with (37) establishes the first claim (i). Similarly, choosing $t=\sigma_{j}^{1-\theta}n^{-1/2+\epsilon/2}$ in the previous argument leads to the second claim (ii).
For the fourth claim (iv), we decompose the sum over $j=1,\dots,p$ along the indices in $J(k_{n})$ and $J(k_{n})^{c}$. To bound the sum over $J(k_{n})^{c}$, we may use (ii) to obtain
$\sum_{j\in J(k_{n})^{c}}\mathbf{P}\Big(|\tilde{X}_{j}-\mu_{j}|\geq C\sigma_{j}^{1-\theta}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ \sum_{j\in J(k_{n})^{c}}(c\sigma_{j})^{\frac{\theta}{c}b_{n}}$
$\ \lesssim\ \sum_{j\geq k_{n}}(Cj^{-\beta})^{\frac{\theta}{c}b_{n}}$
$\ \lesssim\ k_{n}^{{-\frac{\beta\theta}{c}b_{n}}+1}$
$\ \lesssim\ k_{n}^{-\log(n)/c}.$
Regarding the sum over $j\in J(k_{n})$, note that $C\sigma_{j}^{1-\theta}\geq\sigma_{j}$ holds for all $j=1,\dots,p$. Therefore the bound (i) gives
$\sum_{j\in J(k_{n})}\mathbf{P}\Big(|\tilde{X}_{j}-\mu_{j}|\geq C\sigma_{j}^{1-\theta}n^{-1/2+\epsilon/2}\Big)$ $\ \leq\sum_{j\in J(k_{n})}\mathbf{P}\Big(|\tilde{X}_{j}-\mu_{j}|\geq C\sigma_{j}n^{-1/2+\epsilon/2}\Big)$
$\ \lesssim\ k_{n}\Big(\frac{c}{n}\Big)^{\frac{\epsilon}{c}b_{n}}$
$\ \lesssim\ n^{-(2+\delta)}.$
This leads to the third claim (iii) and completes the proof. ∎
Lemma 13.
Fix any constant $\theta\in(0,1)$, and suppose that the conditions of Theorem 1 hold. Then, there is a constant $c\geq 1$ not depending on $n$, such that the following bounds hold for any $j\in\{1,\dots,p\}$,
$\mathbf{P}\Big(|\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}|>\sigma_{j}^{2}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ \Big(\frac{c}{n}\Big)^{\frac{\epsilon}{c}b_{n}}$ (i)
$\mathbf{P}\Big(|\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}|\geq\sigma_{j}^{2-2\theta}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ (c\sigma_{j}^{2})^{\frac{\theta}{c}b_{n}}.$ (ii)
Furthermore, we have
$\sum_{j\in J(k_{n})}\mathbf{P}\Big(|\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}|>C^{2}\sigma_{j}^{2}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ n^{-(2+\delta)}.$ (iii)
$\sum_{j=1}^{p}\mathbf{P}\Big(|\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}|>C^{2}\sigma_{j}^{2-2\theta}n^{-1/2+\epsilon/2}\Big)$ $\ \lesssim\ k_{n}^{-\log(n)/c}.$ (iv)
Proof. The proof of Lemma 12 can be repeated with the i.i.d. random variables $\frac{1}{2}(X_{ij}-X_{i^{\prime}j})^{2}$, playing the role that $X_{ij}$ previously did. Also note that because $\textup{var}\big(\frac{1}{2}(X_{ij}-X_{i^{\prime}j})^{2}\big)\lesssim\sigma_{j}^{4}$ holds under Assumption 1((i)), the parameter $\sigma_{j}^{4}$ plays the role that $\sigma_{j}^{2}$ did in the context of Lemma 12. ∎
For the next lemma, recall that for each $l\in\{1,\dots,b_{n}\}$, the $l$th blockwise variance estimate $\bar{\sigma}_{j}^{2}(l)$ for $\sigma_{j}^{2}$ is defined in (5).
Lemma 14.
Fix any constant $\theta\in(0,1)$, and suppose that the conditions of Theorem 1 hold. Then, there is a constant $c\geq 1$ not depending on $n$, such that the following bound holds for any $j\in\{1,\dots,p\}$ and any $l\in\{1,\dots,b_{n}\}$,
$\mathbf{P}\big(\bar{\sigma}_{j}^{2}(l)\leq\sigma_{j}^{2+2\theta}\big)\ \lesssim\ (c\ell_{n}\sigma_{j}^{2\theta})^{\ell_{n}/4}.$
Proof. Because the $\ell_{n}/2$ terms in the definition of $\bar{\sigma}_{j}^{2}(l)$ are i.i.d., it follows that
$\begin{split}\mathbf{P}\big(\bar{\sigma}_{j}^{2}(l\big)\ \leq\ \sigma_{j}^{2+2\theta})&\ \leq\ \mathbf{P}\Bigg(\max_{\begin{subarray}{c}i,i^{\prime}\in\mathcal{B}_{l}\\ i^{\prime}-i=\ell_{n}/2\end{subarray}}\textstyle\frac{1}{\ell_{n}}(X_{ij}-X_{i^{\prime}j})^{2}\leq\sigma_{j}^{2+2\theta}\Bigg)\\[5.69046pt] &\ =\ \mathbf{P}\bigg(\textstyle\frac{1}{2\sigma_{j}^{2}}(X_{1j}-X_{(\ell_{n}/2+1)j})^{2}\leq\frac{1}{2}\ell_{n}\sigma_{j}^{2\theta}\bigg)^{\ell_{n}/2}.\end{split}$ (38)
Since the independent random variables $X_{1j}/\sigma_{j}$ and $X_{(1+\ell_{n}/2)j}/\sigma_{j}$ have densities whose $L^{\infty}$ norms are $\mathcal{O}(1)$ under Assumption 1((ii)), it follows from Young’s convolution inequality [stein1971introduction, p.178] that the random variable $\frac{1}{\sqrt{2}\sigma_{j}}(X_{1j}-X_{(\ell_{n}/2+1)j})$ also has a density whose $L^{\infty}$ norm is $\mathcal{O}(1)$, and so
$\mathbf{P}\bigg(\frac{1}{2\sigma_{j}^{2}}\big(X_{1j}-X_{(\ell_{n}/2+1)j}\big)^{2}\leq\frac{1}{2}\ell_{n}\sigma_{j}^{2\theta}\bigg)$ $\,\lesssim\,\big(\textstyle\ell_{n}\sigma_{j}^{2\theta}\big)^{1/2}.$
Combining this with (38) completes the proof.∎
Lemma 15.
Fix any constant $\theta\in(0,1)$, and suppose that the conditions of Theorem 1 hold. Then, there is a constant $c\geq 1$ not depending on $n$, such that the event
$\max_{1\leq j\leq p}\frac{\sigma_{j}^{2+2\theta}}{\hat{\sigma}_{j}^{2}}\leq c$
holds with probability at least $1-cn^{-(2+\delta)}$.
Proof. Let $r_{n}=\lceil n^{\epsilon/(\theta\beta)}\wedge p\rceil$. It follows from Lemma 13(i) that there is a constant $c>0$ not depending on $n$ such that the bound
$\max_{j\in J(r_{n})}\frac{\sigma_{j}^{2}}{\hat{\sigma}_{j}^{2}}\leq c$ (39)
holds with probability at least $1-cn^{-(2+\delta)}$. Therefore, a bound of the same form must also hold for $\max_{j\in J(r_{n})}\frac{\sigma_{j}^{2+2\theta}}{\hat{\sigma}_{j}^{2}}$, since $\max_{1\leq j\leq p}\sigma_{j}^{2\theta}\lesssim 1$.
To complete the proof, it remains to handle the maximum of $\sigma_{j}^{2+2\theta}/\hat{\sigma}_{j}^{2}$ over indices $j$ in the complementary set $J(r_{n})^{c}$. Letting $\xi_{jl}=1\{\bar{\sigma}_{j}^{2}(l)\leq\sigma_{j}^{2+2\theta}\}$ for $l=1,\dots,b_{n}$, Kiefer’s inequality (Lemma 22) implies that the following bound holds for any $j\in J(r_{n})$,
$\mathbf{P}\big(\hat{\sigma}_{j}^{2}\leq\sigma_{j}^{2+2\theta}\big)\ \leq\ \mathbf{P}\bigg(\frac{1}{b_{n}}\sum_{l=1}^{b_{n}}\xi_{jl}\geq\frac{1}{2}\bigg)\ \lesssim\ (e\mathbf{E}(\xi_{j1}))^{b_{n}(\frac{1}{2}-\mathbf{E}(\xi_{j1}))^{2}}.$ (40)
Also, Lemma 14 gives
$\mathbf{E}(\xi_{j1})$ $\ \lesssim\ (c\ell_{n}j^{-2\theta\beta})^{\ell_{n}/4}.$
Therefore, combining with with (40), we conclude that
$\sum_{j\in J(r_{n})^{c}}\mathbf{P}(\hat{\sigma}_{j}^{2}\leq\sigma_{j}^{2+\theta})$ $\ \lesssim\ (c\ell_{n})^{\ell_{n}/4}\sum_{j\geq n^{\frac{\epsilon}{\theta\beta}}}(j^{-\theta\beta\ell_{n}/2})^{b_{n}/c}$
$\ \lesssim\ (c\ell_{n})^{\ell_{n}/4}n^{-\frac{n}{2c}+1}$
$\ \lesssim n^{-(2+\delta)},$
where the last step uses $\ell_{n}\asymp n/\log(n)$. Note also that the final bound $n^{-(2+\delta)}$ can be replaced with any fixed positive power of $n^{-1}$, but the current form is all that is needed. ∎
Appendix G Proof of Proposition 1
To ease notation, we let $q=4+\delta$ throughout the proof.
Elliptical case. Suppose $X_{1}$ is a centered elliptical random vector of the form $X_{1}=\eta_{1}\Sigma^{1/2}Z_{1}/\|Z_{1}\|_{2}$, where $Z_{1}$ is a standard Gaussian $p$-dimensional Gaussian vector, and $\eta_{1}$ is independent of $Z_{1}$ with $\mathbf{E}(\eta_{1}^{2})=p$. We first check the $L^{q}$-$L^{2}$ moment equivalence condition (i) with $q=4+\delta$. Letting $w=\Sigma^{1/2}v$ for a generic vector $v\in\mathbb{R}^{p}$, a direct calculation gives
$\|\langle v,X_{1}\rangle\|_{L^{2}}^{2}$ $=\|\eta_{1}\langle w,Z_{1}/\|Z_{1}\|_{2}\rangle\|_{L^{2}}$
$=\|\eta_{1}\|_{L^{2}}^{2}\|w\|_{2}^{2}/p$
$=\|w\|_{2}^{2}.$
Because the distribution of $U_{1}=Z_{1}/\|Z_{1}\|_{2}$ is invariant to orthogonal transformations, it follows that the random variables $\langle w,U_{1}\rangle$ and $\|w\|_{2}\langle e_{1},U_{1}\rangle$ are equal in distribution, where $e_{1}$ is the first standard basis vector. Therefore,
$\|\langle v,X_{1}\rangle\|_{L^{q}}$ $=\|w\|_{2}\|\eta_{1}\|_{L^{q}}\|U_{11}\|_{L^{q}}.$
The quantity $\|U_{11}\|_{L^{q}}$ is at most of order $1/\sqrt{p}$, which can be shown as follows. Due to the independence of $U_{11}$ and $\|Z_{1}\|_{2}$, we have $\|Z_{11}\|_{L^{q}}=\|\|Z_{1}\|_{2}\|_{L^{q}}\|U_{11}\|_{L^{q}}$. Furthermore, Lyapunov’s inequality gives $\|\|Z_{1}\|_{2}\|_{L^{q}}=\|\|Z_{1}\|_{2}^{2}\|_{L^{q/2}}^{1/2}\geq\|\|Z_{1}\|_{2}^{2}\|_{L^{1}}^{1/2}=\sqrt{p}$. Combining the last several steps and the assumption that $\|\eta_{1}\|_{L^{q}}\lesssim\sqrt{p}$, we conclude that
$\|\langle v,X_{1}\rangle\|_{L^{q}}\lesssim\|w\|_{2}=\|\langle v,X_{1}\rangle\|_{L^{2}},$ (41)
which verifies condition (i).
Regarding the density condition (ii), note that if $e_{j}$ is the $j$th standard basis vector, then the vector $w=\Sigma^{1/2}e_{j}/\sigma_{j}$ satisfies $\|w\|_{2}=1$. So, the discussion above shows that all the random variables $X_{11}/\sigma_{1},\dots,X_{1p}/\sigma_{p}$ have the same distribution, which is that of $\eta_{1}\langle e_{1},U_{1}\rangle$. In particular, if the random variable $X_{11}/\sigma_{1}$ has a Lebesgue density $f_{1}$ such that $\|f_{1}\|_{L^{\infty}}\lesssim 1$, then condition (ii) holds, which completes the proof in the elliptical case.
Separable case. Suppose that $X_{1}$ has a centered separable distribution so that $X_{1}=\Sigma^{1/2}\zeta_{1}$, where $\zeta_{1}=(\zeta_{11},\dots,\zeta_{1p})$ has i.i.d. entries with $\mathbf{E}(\zeta_{11})=0$, and $\textup{var}(\zeta_{11})=1$. To check the $L^{q}$-$L^{2}$ moment equivalence condition (i), it suffices to show that $\|\langle w,\zeta_{1}\rangle\|_{L^{q}}\lesssim\|\langle w,\zeta_{1}\rangle\|_{L^{2}}$ for any $w\in\mathbb{R}^{p}$. Using Rosenthal’s inequality (Lemma 16), and the assumption that $\max_{1\leq j\leq p}\|\zeta_{1j}\|_{L^{q}}\lesssim 1$, we have
$\|\langle w,\zeta_{1}\big\rangle\|_{L^{q}}$ $\lesssim\max\bigg\{\|\langle w,\zeta_{1}\rangle\big\|_{L^{2}},\Big(\sum_{j=1}^{p}\big\|w_{j}\zeta_{1j}\|_{L^{q}}^{q}\Big)^{1/q}\bigg\}$
$\lesssim\max\big\{\|w\|_{2},\|w\|_{q}\big\}$
$=\|\langle w,\zeta_{1}\rangle\|_{L^{2}}$
as needed. Finally, the density condition (ii) is a direct consequence of Theorem 1.2 in the paper [rudelson2015small] and the assumption that $\max_{1\leq j\leq p}\|g_{j}\|_{L^{\infty}}\lesssim 1$, where $g_{j}$ is the Lebesgue density of $\zeta_{1j}$. ∎
Appendix H Background results
Lemma 16 (Rosenthal inequalities [Rosenthal]).
Fix $q\geq 1$, and let $\xi_{1},\ldots,\xi_{n}$ be independent random variables. Then, there is an absolute constant $c>0$ such that the following two statements are true.
(i). If $\xi_{1},\ldots,\xi_{n}$ are non-negative, then
$\Big\|\sum_{i=1}^{n}\xi_{i}\Big\|_{L^{q}}\leq c\cdot q\cdot\max\bigg\{\Big\|\sum_{i=1}^{n}\xi_{i}\Big\|_{L^{1}},\Big(\sum_{i=1}^{n}\big\|\xi_{i}\big\|_{L^{q}}^{q}\Big)^{1/q}\bigg\}.$
(ii). If $q\geq 2$, and $\xi_{1},\ldots,\xi_{n}$ are centered, then
$\Big\|\sum_{i=1}^{n}\xi_{i}\Big\|_{L^{q}}\leq c\cdot q\cdot\max\bigg\{\Big\|\sum_{i=1}^{n}\xi_{i}\Big\|_{L^{2}},\Big(\sum_{i=1}^{n}\big\|\xi_{i}\big\|_{L^{q}}^{q}\Big)^{1/q}\bigg\}.$
For the next lemma, recall that we denote the stable rank of a non-zero positive semidefinite matrix $A$ as ${\tt{r}}(A)=\textup{tr}(A)^{2}/\|A\|_{F}^{2}$.
Lemma 17 (Lower-tail bound for Gaussian maxima [lopes2022sharp]).
Let $\xi\sim\mathcal{N}(0,R)$ be a Gaussian random vector in $\mathbb{R}^{d}$ for some correlation matrix $R$, and fix two constants $a,b\in(0,1)$ with respect to $d$. Then, there is a constant $c>0$ depending only on $(a,b)$ such that the following inequality holds for any integer $k$ satisfying $2\leq k\leq\frac{b^{2}}{4}{\tt{r}}(R)$,
$\mathbf{P}\Big(\max_{1\leq j\leq d}\xi_{j}\leq a\sqrt{2(1-b)\log(k)}\Big)\leq c\,k^{\frac{-(1-b)(1-a)^{2}}{b}}\big(\log(k)\big)^{\frac{1-b(2-a)-a}{2b}}$
The next result is a variant of Lemma A.7 in [spokoiny2015bootstrap] that can be proven in essentially the same way.
Lemma 18 (Gaussian comparison inequality).
Let $\zeta\sim N(0,I_{d})$ and $\xi\sim N(0,A)$ for some positive semidefinite matrix $A\in\mathbb{R}^{d\times d}$. Then,
$\sup_{s\in\mathbb{R}}\bigg|\mathbf{P}\Big(\max_{1\leq j\leq d}\xi_{j}\leq s\Big)-\mathbf{P}\Big(\max_{1\leq j\leq d}\zeta_{j}\leq s\Big)\bigg|\ \leq\ 2\|A-I_{d}\|_{F}.$
Lemma 19 (Bentkus’ Berry-Esseen Theorem [Bentkus:2003]).
Let $V_{1},\ldots,V_{n}$ be i.i.d. random vectors in $\mathbb{R}^{d}$ with zero mean and identity covariance matrix. Furthermore, let $\gamma_{d}$ denote the standard Gaussian distribution on $\mathbb{R}^{d}$, and let $\mathscr{A}$ denote the collection of all Borel convex subsets of $\mathbb{R}^{d}$. Then, there is an absolute constant $c>0$ such that
$\sup_{\mathcal{A}\in\mathscr{A}}\left|\mathbf{P}\Big(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}V_{i}\in\mathcal{A}\Big)-\gamma_{d}(\mathcal{A})\right|\leq\frac{c\cdot d^{1/4}\cdot\mathbf{E}\left\|V_{1}\right\|_{2}^{3}}{n^{1/2}}.$
Lemma 20 (Nazarov’s inequality).
Let $(\zeta_{1},\ldots,\zeta_{d})$ be a Gaussian random vector, and suppose the parameter $\underline{\sigma}^{2}\mathrel{\mathop{\ordinarycolon}}=\min_{1\leq j\leq d}\textup{var}(\zeta_{j})$ is positive. Then, for any fixed $\epsilon>0$,
$\sup_{s\in\mathbb{R}}\mathbf{P}\Big(\big|\max_{1\leq j\leq d}\zeta_{j}-s\big|\leq\epsilon\big)\ \leq\ \frac{2\epsilon}{\underline{\sigma}}(\sqrt{2\log(d)}+2).$
This version of Nazarov’s inequality appears in Lemma 4.3 of [chernozhukov2016empirical] and originates from [nazarov2003maximal].
Lemma 21 (Fuk-Nagaev inequality).
Fix $q\geq 1$, and let $\xi_{1},\ldots,\xi_{n}$ be centered independent random variables. Then, for any fixed $s>0$,
$\mathbf{P}\bigg(\Big|\sum_{i=1}^{n}\xi_{i}\Big|\geq s\bigg)\ \leq\ 2\Big(\frac{q+2}{qs}\Big)^{q}\sum_{i=1}^{n}\mathbf{E}|\xi_{i}|^{q}+2\exp\bigg(\frac{-2s^{2}}{(q+2)^{2}e^{q}\sum_{i=1}^{n}\mathbf{E}(\xi_{i}^{2})}\bigg).$
This statement of the Fuk-Nagaev inequality is based on [rio2017constants, eqn. 1.7].
Lemma 22 (Kiefer’s inequality).
If $\xi_{1},\ldots,\xi_{n}$ are i.i.d. Bernoulli random variables, then
$\mathbf{P}\bigg(\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\geq\frac{1}{2}\Big)\ \leq\ 2\big(e\mathbf{E}(\xi_{1})\big)^{n(1/2-\mathbf{E}(\xi_{1}))^{2}}.$
The result above is the modification of Corollary A.6.3 in [van1996weak]. In that reference, the success probability $\mathbf{E}(\xi_{1})$ of the Bernoulli random variables is assumed to be less than $1/e$, but in the formulation above, the result holds for all values of $\mathbf{E}(\xi_{1})$.